### v2

# Operational reconstruction of Feynman rules for quantum amplitudes via composition algebras

Abstract
This article explores an operational model for transition amplitudes between measurements proposed by Goyal et al. within the quantum reconstruction program. To classify suitable amplitude algebras, we distinguish mathematical axioms, physical choices, and their consequences. This leads to several improvements on the published work: Our coordinate-independent approach requires no two-dimensional amplitudes a priori. All scalar field and vector space axioms are traced from model axioms and observer choices, including additive and multiplicative units and inverses. Existing mathematical characterizations identify allowable amplitude algebras as the real associative composition algebras, namely the complex numbers and the quaternions, as well as their split forms. Observed probabilities are quadratic in amplitudes, akin to the Born rule. We examine selected implications of the proposed axioms, reformulate observer questions, and highlight the broad applicability of our framework to subsequent discovery.
Addresses jenskoeplinger@gmail.com, michael.habeck@uni-jena.de, pgoyal@albany.edu
1 Introduction
1.1 Context
The Dirac–von Neumann axioms of quantum theory possess a rich mathematical structure. However, unlike the theories of physics which preceded it (such as Newtonian mechanics and Maxwell’s theory of electromagnetism), it is difficult to account for this structure by pointing to specific physical principles, facts about the physical world, or theoretical choices in how we represent that world symbolically. This predicament is the legacy of the rather convoluted process by which the quantum formalism was historically obtained, a process that heavily relied on inspired mathematical guesswork, as well as new physical principles and phenomena.
Since the creation of quantum formalism a century ago, there have been many attempts to rectify this situation (e.g. [BvN1936, Mackey1957, EP1963, Fivel1994]). However, these attempts have had comparatively little impact, in part due to the rather abstract nature of some of their assumptions. The rise of quantum information in the 1980s led to the hypothesis that quantum theory might be derivable in terms of information-theoretic principles [Rovelli1996, Fuchs2003], and to the conviction such a derivation could be achieved without recourse to abstract mathematical assumptions that lack clear physical justification [Hardy2001, Grin2003, Grin2007, Goyal2023]. Following promising early information-inspired reconstructive attempts (e.g. [Wootters1980]), one of the first successful reconstructions of the finite-dimensional quantum axioms was carried out by Lucien Hardy [Hardy2001, Hardy2001whyQ] in 2001. In the subsequent years, numerous other reconstructions of the finite-dimensional axioms followed (e.g. [DAriano2007, Goyal2008, BD2010, GKS2010, CPDA2011, MMAP2013, Goyal2014, MM2016, Hoehn2017, SSC2018]), which approached the reconstructive challenge using a wide variety of mathematical and conceptual tools.
One of the desired characteristics of a reconstruction is that it enables one to clearly grasp why the complex number system is appropriate for the description of quantum phenomena. In this regard, reconstructions of the Feynman formulation of quantum theory [Feynman1948, FeH1965] – as opposed to the standard finite-dimensional Dirac–von Neumann axioms – are of special interest. The standard formalism attributes quantum states to physical systems, stipulates how these evolve and update in response to interactions with, and measurements upon, the system. In contrast, the Feynman formulation eschews the notion of quantum state and instead ascribes complex-valued amplitudes to paths that a system may take between detection events. The resulting formalism is dramatically simpler than the standard state-based formalism. Its three core rules prescribe how to combine these amplitudes according to whether these paths are in series (1st rule) or in parallel (2nd), and how to compute transition probabilities from these amplitudes (3rd). Additional rules show, for example, how to combine amplitudes belonging to transitions of non-interacting systems (which corresponds to the usual tensor product rule in the Dirac–von Neumann axioms), and how to handle identical particle systems.
As shown in [Tiko1988, Caticha1998, Tiko1999, GKS2010], Feynman’s core rules possess striking symmetries (such as associativity, commutativity, distributivity, and a product constraint for probabilities). Furthermore, as shown in [Caticha1998, GKS2010], these symmetries can be justified largely on operational grounds, without explicit reference to specific quantum phenomena, and moreover can be harnessed to reconstruct Feynman’s rules. However, these reconstructions leave something to be desired. For example, at the outset, [GKS2010] posits that each path is represented by a pair of real numbers, thereby ruling out algebraic possibilities beyond the complex numbers a priori. It also employs a rather ad hoc approach to the solution of the symmetry equations, which leads to additional hitherto unresolved issues [Koepl2023gks].
Our goal in the current paper is to harness the above-mentioned operationally-established symmetries by determining a direct connection between these symmetries and known foundational theorems in mathematics. There is rich historical context in linear algebra, from associative division algebras (the reals, complexes, and quaternions [Frob1878]), multiplicative sums of squares (also permitting nonassociative octonions [Hurwitz1898]), to a complete classification of algebras permitting a quadratic form $p$ with product composition, $p\left(ab\right)=p\left(a\right)p\left(b\right)$ (composition algebras; [Jacobson1958, Jacobson1981]). Our approach rests on the observation that there is substantial overlap between the symmetries and structures formulated in [GKS2010, Goyal2014] and the underlying axioms of the theorems that characterize composition algebras. By judiciously expanding the set of operations formulated in [GKS2010], and placing emphasis on specific types of paths, we are able to show that complex numbers together with a small number of other possible algebras are suitable to represent the operational model [^1] This in turn promises advancements for an existing body of related research [BCL1990, Goyal2012, Goyal2014, Goyal2015], and may provide a clearer understanding of alternative postulates introduced in [GK2011, SK2019], some of which faced critiques for their perceived unnecessary strength (e.g. [HSP2014])..
The remainder of the paper is organized as follows. In section 2, we recap the original model from [GKS2010] and develop select algebraic properties. We avoid complete classification of this model algebra, to not attempt to optimize what is supposed to be a building block towards an unknown future completion. The question at hand is then specified in section 3, asking for probabilities of possible outcomes given an experimental setup. To answer this question, we introduce the choice representation of amplitudes in section 4, supported by a simplicity argument towards answering the question posed. Finally, in section 5, we demonstrate that allowable probabilities are quadratic in amplitudes, similar to the Born rule from canonical physics, and classify allowable amplitude algebras as the associative composition algebras over the reals. We close with an outlook (section 6) that traces what was accomplished, points out nontrivial consequences of seemingly minute tweaks to the model, and advertises exploration of one specific generalization. Algebras that become permissible under this abstract generalization match those from contemporary algebraic representation of observed symmetries and structure of the fundamental particles and forces in nature [Furey2023rmap1, FuH2024]. This is surprising – to say the least – given the different provenance. Rather than taking quantum theory as a given, we invite to look at it from a new angle altogether.
1.2 Methodology
The methodology in this paper is structured along three primary axes: axiomatic definition, choice postulates, and derivation validation:
  • Axioms provide the foundational assumptions that determine the model’s relevance and scope.
  • Choices are further postulates based on specific questions asked from the model. They provide the template into which the inherent properties of the phenomenon under study are to express themselves.
  • Derivations then expose true statements that follow from axioms and choices through deductive reasoning. This provides the basis for evaluating choices for their logical coherence, to support their claimed generality, and to ensure they do not conflict with the model.
Validation of axioms and choices ultimately resides in experimental observation: Good choices yield useful answers, given the questions asked. While mathematically, the difference between axioms and choices may be little to none, the separation here aims at usefulness in physics:
  • From the ground up: The axiomatic model is advertised as a primitive building block in a way that promises broad utility towards a future complete reconstruction of Quantum Mechanics. Clarity in axioms and subsequent choices aims at providing researchers with an easier selection on what to carry forward, what to modify, and what consequences may follow.
  • With hindsight from canonical physics: After postulating a choice representation of the model in terms of amplitudes, we will arrive at a finite set of allowable amplitude algebras, which includes the complex numbers. Any faithful reconstruction of canonical Quantum Mechanics must explain the utility of the complexes. By highlighting the choices we are making in representing our model algebra, tailored to answer a specific question, we aim to clarify how to arrive at the complexes.
2 Model
We adopt the operational model from [GKS2010], which constructs experiments on quantum systems by arranging several measurements in sequence. Each measurement comprises detectors, which are distinct and non-filtering: Exactly one detector fires in each measurement, corresponding to a measurement result. A sequence of measurement results is then called a path. Sequential composition and simultaneous merging of paths yield a model algebra with basic operations chaining and coarsening, respectively. In this section, we develop the properties and relations of this model algebra [^2] The term “algebra” here is understood in the sense of universal algebra as a space where certain operations act on some set of elements. It is the “logic” of the model per [GKS2010]., including inversion, reversal, and insertion.
Interactions between different quantum systems are assumed to be happening between measurements. Together with spacetime and dynamics, interactions are out of scope here.
2.1 Measurements with possible measurement outcomes
We denote an experiment as a sequence $\mathcal{M}$ with $\ell\geq 2$ measurements $\mathbf{M}_{j}$, written
$\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right].$ (2.1)
The source $\mathbf{S}$ is the first measurement in a sequence, $\mathbf{S}\left(\mathcal{M}\right)=\mathbf{M}_{1}$, and the target $\mathbf{T}$ the last, $\mathbf{T}\left(\mathcal{M}\right)=\mathbf{M}_{\ell}$. Each measurement comprises a finite number of detectors, $\mathbf{M}_{j}=\left\{M_{j}^{1},M_{j}^{2},M_{j}^{3},\ldots\right\}$, which form the set of possible outcomes for that measurement.
There are two classes of detectors $M_{j}^{r}$, atomic and non-atomic. If all detectors in a measurement are atomic, we call the measurement itself atomic. By postulate, we require the source and target of a sequence to be atomic.
Two separate measurements are equal, $\mathbf{M}_{j}=\mathbf{M}_{k}$, if they have identical possible outcomes and with that, are modeled with the same set of detectors. A sequence that consists only of equal measurements is called trivial.
We write a measurement’s underlying atomic outcome elements as $m,m^{\prime},m^{\prime\prime},\ldots$. Actual outcomes then are from a partition over these underlying elements. For example, atomic outcomes are sets with one element, $\left\{m\right\}$, $\left\{m^{\prime}\right\}$, $\left\{m^{\prime\prime}\right\}$, …, and coarse-grained outcomes are sets with two or more elements, $\left\{m,m^{\prime}\right\}$, $\left\{m,m^{\prime\prime}\right\}$, $\left\{m,m^{\prime},m^{\prime\prime}\right\}$, …
We use raised-prefix notation $^{\alpha}\mathbf{M},{}^{\beta}\mathbf{M},{}^{\gamma}\mathbf{M},\ldots,{}^{\cup}\mathbf{M}$ to label different partitions over the same underlying element set and call these measurements weakly equivalent, $^{\alpha}\mathbf{M}\sim{}^{\beta}\mathbf{M}\sim{}^{\gamma}\mathbf{M}\sim\ldots\sim{}^{\cup}\mathbf{M}$. Prefix $\alpha$ is reserved for atomic measurements, and prefix $\cup$ for measurements that are fully coarse-grained, i.e., consisting of a single detector with an outcome that is the set of all elements.
For example, five possible measurements can be built over the set with three underlying atomic outcome elements $m,m^{\prime},m^{\prime\prime}$, corresponding to the five possible partitions of a set with three elements. These measurements are weakly equivalent (see Figure 2.1):
Figure 2.1: Given three underlying atomic outcome elements $m,m^{\prime},m^{\prime\prime}$, the five possible partitions are the measurements $^{\alpha}\mathbf{M},{}^{\beta}\mathbf{M},{}^{\gamma}\mathbf{M},{}^{\delta}\mathbf{M},{}^{\cup}\mathbf{M}$. The elements in each partition are its detectors.
$^{\alpha}\mathbf{M}$ $=\left\{\left\{m\right\},\left\{m^{\prime}\right\},\left\{m^{\prime\prime}\right\}\right\}$
$^{\beta}\mathbf{M}$ $=\left\{\left\{m,m^{\prime}\right\},\left\{m^{\prime\prime}\right\}\right\}$
$^{\gamma}\mathbf{M}$ $=\left\{\left\{m\right\},\left\{m^{\prime},m^{\prime\prime}\right\}\right\}$ (2.2)
$^{\delta}\mathbf{M}$ $=\left\{\left\{m,m^{\prime\prime}\right\},\left\{m^{\prime}\right\}\right\}$
$^{\cup}\mathbf{M}$ $=\left\{\left\{m,m^{\prime},m^{\prime\prime}\right\}\right\}$
$^{\alpha}\mathbf{M}$ $\sim{}^{\beta}\mathbf{M}\sim{}^{\gamma}\mathbf{M}\sim{}^{\delta}\mathbf{M}\sim{}^{\cup}\mathbf{M}.$ (2.3)
2.2 Paths with actual measurement results
Given a measurement sequence $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]$, paths $A,B,C,\ldots$ are sequences of measurement results from the possible detectors at each measurement. The set of all allowable paths over $\mathcal{M}$ is written as $\mathcal{S_{M}}$.
Given a path $A\in\mathcal{S_{M}}$, we write $A_{j}$ for the detector result at measurement $\mathbf{M}_{j}$, i.e., $A_{j}\in\mathbf{M}_{j}$,
$A=\left[A_{1},A_{2},\ldots,A_{\ell}\right].$ (2.4)
Just as the source and target in a path are the respective measurements in a sequence, $\mathbf{S}\left(A\right)=\mathbf{S}\left(\mathcal{M}\right)$ and $\mathbf{T}\left(A\right)=\mathbf{T}\left(\mathcal{M}\right)$, we call the first and last measurement outcome in a path the source result $S\left(A\right)=A_{1}$ and target result $T\left(A\right)=A_{\ell}$. Because the source and target of a measurement sequence must be atomic, this will be the case for source and target results.
2.2.1 Preparation and selection
An experiment always consists of at least two measurements ($\ell\geq 2$). The first measurement, $\mathbf{M}_{1}$, is required to be atomic and prepares the system such that the experiment is well-defined and can be repeated. The last measurement, $\mathbf{M}_{\ell}$, is also atomic and yields a specific outcome. Therefore, measurement sequences $\left[\mathbf{M}_{1},\mathbf{M}_{2}\right]$ of length $\ell=2$, with paths $\left[M_{1},M_{2}\right]$, are the shortest possible sequences.
2.2.2 Cyclic, symmetric (palindromic), and trivial paths
A path $^{\circ}\negthinspace A$ is cyclic if its source and target are identical, i.e., $\mathbf{M}_{1}=\mathbf{M}_{\ell}$ with $S\left({}^{\circ}\negthinspace A\right)=T\left({}^{\circ}\negthinspace A\right)$. Furthermore, cyclic paths are symmetric (or palindromic), $^{\circledcirc}\negthinspace A$, if their measurements and outcomes are identical in the opposite order: $\mathbf{M}_{1}=\mathbf{M}_{\ell}$ with result $^{\circledcirc}\negthinspace A_{1}=^{\circledcirc}\negthinspace A_{\ell}$, $\mathbf{M}_{2}=\mathbf{M}_{\ell-1}$ with $^{\circledcirc}\negthinspace A_{2}=^{\circledcirc}\negthinspace A_{\ell-1}$, and so on. Note that the definitions for cyclic and symmetric paths require identical outcomes and also equal measurements at the relevant sequence positions: Paths over cyclic (or symmetric) measurement sequences generally do not have cyclic (or symmetric) outcomes, and conversely, a path with measurement outcomes matching these patterns does not require matching measurement patterns due to the weak equivalence freedom for measurements.
Take, for example, the two measurements $^{\alpha}\mathbf{M}=\left\{\left\{m\right\},\left\{m^{\prime}\right\},\left\{m^{\prime\prime}\right\}\right\}$ and $^{\delta}\mathbf{M}=\left\{\left\{m,m^{\prime\prime}\right\},\left\{m^{\prime}\right\}\right\}$ from equation (2.2) above. Both have detector $M^{\prime}=\left\{m^{\prime}\right\}$ as a possible outcome. The $^{\alpha}\mathbf{M}$ and $^{\delta}\mathbf{M}$ are weakly equivalent but not equal. Now take two more arbitrary measurements, $\mathbf{N}$ and $\mathbf{O}$, with results $N\in\mathbf{N}$ and $O\in\mathbf{O}$, respectively, and form a measurement sequence $\left[\mathbf{N},{}^{\alpha}\mathbf{M},\mathbf{O},{}^{\delta}\mathbf{M},\mathbf{N}\right]$. This sequence allows for results $\left[N,M^{\prime},O,M^{\prime},N\right]$, which would be considered cyclic but not symmetric because $^{\alpha}\mathbf{M}\neq{}^{\delta}\mathbf{M}$.
Finally, a trivial path consists only of identical measurement results over equal measurements.
2.3 Model algebra
2.3.1 Chaining and coarsening
We follow the original approach in [GKS2010] and develop algebraic relations for measurement sequences and paths. In order to avoid confusion with established mathematical terms like “sequence algebra” or “path algebra,” we adopt a generic term model algebra here. It simply denotes algebraic relations that trace from the operational model in this paper without attempting to specify their general context [^3] That is not to say that exploring the model algebra by itself in the context of established mathematical frameworks would not be interesting. Doing so is bound to provide fruitful insight. As we make certain choices in this paper, we aim to allow for the widest possible interpretation and generalization of our results. Narrowing the model algebra at this point is unneeded and, therefore, unwanted for our purpose..
The space of all measurements is now written as $\mathbb{M}$. Experiments then are from the possible sequences of measurements, written $\left[\mathbb{M}_{j}\right]_{j\geq 2}$, or simply $\left[\mathbb{M}\right]$. Two measurement sequences $\mathcal{M}^{\mathrm{A}}=\left[\mathbf{M}_{1}^{\mathrm{A}},\mathbf{M}_{2}^{\mathrm{A}},\ldots,\mathbf{M}_{\ell_{\mathrm{A}}}^{\mathrm{A}}\right]$ and $\mathcal{M}^{\mathrm{B}}=\left[\mathbf{M}_{1}^{\mathrm{B}},\mathbf{M}_{2}^{\mathrm{B}},\ldots,\mathbf{M}_{\ell_{\mathrm{B}}}^{\mathrm{B}}\right]$, $\mathcal{M}^{\mathrm{A}},\mathcal{M}^{\mathrm{B}}\in\left[\mathbb{M}\right]$, where the target of $\mathcal{M}^{\mathrm{A}}$ is equal to the source of $\mathcal{M}^{\mathrm{B}}$, may be composed sequentially to form a new sequence $\mathcal{M}^{\mathrm{C}}$:
$\cdot$ $:\left[\mathbb{M}\right]\times\left[\mathbb{M}\right]\rightharpoonup\left[\mathbb{M}\right],$
$\mathcal{M}^{\mathrm{A}},\mathcal{M}^{\mathrm{B}},\mathcal{M}^{\mathrm{C}}$ $\in\left[\mathbb{M}\right],$ (2.5)
$\textrm{with }\mathbf{T}\left(\mathcal{M^{\mathrm{A}}}\right)=\mathbf{S}\left(\mathcal{M}^{\mathrm{B}}\right),$
$\mathcal{M}^{\mathrm{C}}=\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$ $:=\left[\mathbf{M}_{1}^{\mathrm{A}},\ldots,\mathbf{M}_{\ell_{\mathrm{A}}}^{\mathrm{A}}=\mathbf{M}_{1}^{\mathrm{B}},\ldots,\mathbf{M}_{\ell_{\mathrm{B}}}^{\mathrm{B}}\right],$
$\ell_{\mathrm{C}}$ $=\ell_{\mathrm{A}}+\ell_{\mathrm{B}}-1.$ (2.6)
[…]
If paths $A,B$ exist over these measurement sequences, $A\in\mathcal{S}_{\mathcal{M}^{\mathrm{A}}}$, $B\in\mathcal{S}_{\mathcal{M}^{\mathrm{B}}}$, where likewise the target of $A$ matches the source of $B$, then the same operation is defined for measurement paths as well:
$\cdot$ $:\mathcal{S}_{\left[\mathbb{M}\right]}\times\mathcal{S}_{\left[\mathbb{M}\right]}\rightharpoonup\mathcal{S}_{\left[\mathbb{M}\right]},$
$A$ $\in\mathcal{S}_{\mathcal{M}^{\mathrm{A}}},$
$B$ $\in\mathcal{S}_{\mathcal{M}^{\mathrm{B}}},$ (2.7)
$\begin{array}[]{rl}\textrm{with}&\mathbf{T}\left(\mathcal{M^{\mathrm{A}}}\right)=\mathbf{S}\left(\mathcal{M}^{\mathrm{B}}\right),\\ &\mathbf{t}\left(A\right)=\mathbf{s}\left(B\right),\end{array}$ (2.10)
$A\cdot B$ $=\left[A_{1},A_{2},\ldots,A_{\ell_{A}}=B_{1},B_{2},\ldots,B_{\ell_{B}}\right],$
$\ell_{A\cdot B}$ $=\ell_{A}+\ell_{B}-1.$ (2.11)
Figure 2.2: Chaining ($\cdot$) and coarsening ($\vee$) of sequences of measurements (experiments; $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$) and measurement results (paths; $A\cdot B$, $A\vee B$).
Another operation on paths is coarsening, which is a form of simultaneous merging of possible detector outcomes and measurement results on a path. Given a measurement sequence $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]$, coarsening modifies exactly one of these $\mathbf{M}_{j}\in\mathcal{M}$ that is neither source or target [^4] Source and target must be atomic and, therefore, could not result in an allowable path if either were to be coarsened., $1<j<\ell$, by replacing it with a coarsened version, $^{\vee}\mathbf{M}_{j}\in\mathbb{M}$. All detector outcomes of $\mathbf{M}_{j}$ are subsets of the outcomes of the newly formed, different measurement $^{\vee}\mathbf{M}_{j}$ of lesser cardinality, $\left|{}^{\vee}\mathbf{M}_{j}\right|<\left|\mathbf{M}_{j}\right|$, and the $^{\vee}\mathbf{M}_{j}$ are always non-atomic. This requires $\mathbf{M}_{j}$ and $^{\vee}\mathbf{M}_{j}$ to be weakly equivalent but not equal:
$\mathbf{M}_{j}\sim{}^{\vee}\mathbf{M}_{j},\,\,\,\,\,\,\,\,\mathbf{M}_{j}\neq{}^{\vee}\mathbf{M}_{j},\,\,\,\,\,\,\,\,\left|{}^{\vee}\mathbf{M}_{j}\right|<\left|\mathbf{M}_{j}\right|.$ (2.12)
We write $\mathcal{M}^{\vee\left\{j\right\}}$ for the sequence that differs from $\mathcal{M}$ exactly by replacing the measurement $\mathbf{M}_{j}$ with $^{\vee}\mathbf{M}_{j}$.
Formally, take two paths $A,B\in\mathcal{S}_{\mathcal{M}}$ that differ only at $\mathbf{M}_{j}$ by $A_{j},B_{j}\in\mathbf{M}_{j}$, $A_{j}\in A$, $B_{j}\in B$, $A_{j}\cap B_{j}=\emptyset$, and form the new measurement $^{\vee}\mathbf{M}_{j}$ by coarsening these two detector outcomes to become a single new coarse-grained outcome $A_{j}\cup B_{j}\in{}^{\vee}\mathbf{M}_{j}$. Coarsening then yields a path that contains this coarsened outcome as a result:
$\vee$ $:\mathcal{S}_{\left[\mathbb{M}\right]}\times\mathcal{S}_{\left[\mathbb{M}\right]}\rightharpoonup\mathcal{S}_{\left[\mathbb{M}\right]},$
$A,B$ $\in\mathcal{S}_{\mathcal{M}},$
$C$ $\in\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}},$ (2.13)
$\begin{array}[]{rl}\textrm{with}&1<j<\ell_{A}=\ell_{B},\\ &A_{j},B_{j}\in\mathbf{M}_{j}\in\mathcal{M},\\ &A_{j}\cap B_{j}=\emptyset,\\ &C_{k}:=A_{k}=B_{k}\,\left(\forall k\neq j\right),\\ &C_{j}:=A_{j}\cup B_{j}\in{}^{\vee}\mathbf{M}_{j}\in\mathcal{M}^{\lor\left\{j\right\}},\\ &\left|{}^{\vee}\mathbf{M}_{j}\right|=\left|\mathbf{M}_{j}\right|-1,\end{array}$ (2.20)
$A\vee B$ $:=C.$
Similar to chaining, coarsening $\vee:\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\times\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\rightharpoonup\mathcal{S}_{\left[\mathbb{M}\right]}$ is a partial binary operation, i.e., it is only defined on a subset of compatible elements from $\mathcal{S}_{\left[\mathbb{M}\right]}$. See Figure 2.2 for an illustration.
2.3.2 Chaining is associative but partially noncommutative
Sequences of measurements and outcomes are strictly ordered. The same experiment can be conceptually divided in different ways. When chaining three or more sequences, which sequence pairs are chained first is irrelevant: The operation is associative. For suitable $\mathcal{M}^{\mathrm{A}},\mathcal{M}^{\mathrm{B}},\mathcal{M}^{\mathrm{C}}\in\mathbb{\left[\mathbb{M}\right]}$, $A\in\mathcal{S}_{\mathcal{M}^{\mathrm{A}}}$, $B\in\mathcal{S}_{\mathcal{M}^{\mathrm{B}}}$, $C\in\mathcal{S}_{\mathcal{M}^{\mathrm{C}}}$, there is
$\left(\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}\right)\cdot\mathcal{M}^{\mathrm{C}}$ $=\mathcal{M}^{\mathrm{A}}\cdot\left(\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{C}}\right),$ (2.21)
$\left(A\cdot B\right)\cdot C$ $=A\cdot\left(B\cdot C\right).$ (2.22)
However, chaining is generally not commutative. This is trivially the case when only one of the commuted expressions for measurement sequences $\left\{\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}},\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{A}}\right\}$ or paths $\left\{A\cdot B,B\cdot A\right\}$ is defined: Only because the target of a $\mathcal{M}^{\mathrm{A}}$ matches the source of $\mathcal{M}^{\mathrm{B}}$, permitting chaining $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$, this does not have to be true when commuting the two, permitting $\mathcal{M}^{\mathrm{B}}\overset{?}{\cdot}\mathcal{M}^{\mathrm{A}}$ in principle.
In the other case, when both commuted expressions $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{A}}$ are defined, there is still no a priori basis for this symmetry. We don’t need to assume that changing the order of measurements or changing the order of measurement outcomes never affects the observation.
Since not every commuted expression is allowable, we call this property of chaining partial noncommutativity: Where defined, chaining may or may not be commutative.
2.3.3 Coarsening is commutative and partially associative
Coarsening acts on the set of possible measurement outcomes by merging two detectors. When a measurement is performed, the corresponding merged results in a path are yielded. Since there is no ordering of detectors that make up a measurement, the coarsening operation is trivially commutative: For suitable $\mathcal{M}\in\mathbb{\left[\mathbb{M}\right]}$, $A,B,C\in\mathcal{S}_{\mathcal{M}}$, there is
$A\vee B=B\vee A.$ (2.23)
Where allowable, coarsening is also associative:
$\left(A\vee B\right)\vee C=A\vee\left(B\vee C\right).$ (2.24)
To account that either side of this associativity relation may be undefined, we call this property partial associativity: Where defined, coarsening is always associative. For both sides of the associativity relation to be defined, the three paths $A,B,C$ must differ exactly (and only) at the same sequence index $j$, i.e., $\left(A\vee B\right)\in\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}}$ and $\left(B\vee C\right)\in\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}}$.
2.3.4 Chaining distributes over coarsening
Figure 2.3: Chaining distributes over coarsening
It can be readily seen that chaining is distributive over coarsening (figure 2.3): For $\mathcal{M}^{\mathrm{A}},\mathcal{M}^{\mathrm{BC}}\in\mathbb{\left[\mathbb{M}\right]}$, $A\in\mathcal{S}_{\mathcal{M}^{\mathrm{A}}}$, $B,C\in\mathcal{S}_{\mathcal{M}^{\mathrm{BC}}}$, there is
$A\cdot\left(B\vee C\right)=A\cdot B\vee A\cdot C.$ (2.25)
The left-hand side of the left-distributivity relation chains a path over a measurement sequence $\mathcal{M}^{\mathrm{A}}$ to a path over $\mathcal{M}^{\mathrm{BC}\lor\left\{j\right\}}$, i.e., over a sequence of measurements that differs from the original $\mathcal{M}^{\mathrm{BC}}$ only at an index $j\neq 1,\ell$. With that, the source of $B,C$ is the same as that of $B\vee C$. Conversely, the right-hand side of the left-distributivity relation extends $B$ and $C$ by an identical path $A$ that is not subject to coarsening.
A similar argument applies to right-distributivity:
$\left(B\vee C\right)\cdot A=B\cdot A\vee C\cdot A.$ (2.26)
2.3.5 Reversal
Reversal takes a sequence of measurements $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]\in\mathbb{\left[\mathbb{M}\right]}$ or measurement results $A=\left[A_{1},A_{2},\ldots,A_{\ell}\right]\in\mathcal{S}_{\mathcal{M}}$ and puts all elements in the opposite order:
[…]
$\overline{\mathcal{M}}$ $=\left[\mathbf{M}_{\ell},\mathbf{M}_{\ell-1},\ldots,\mathbf{M}_{1}\right],$ (2.27)
[…]
$\overline{A}$ $=\left[A_{\ell},A_{\ell-1},\ldots,A_{1}\right].$ (2.28)
This operation replaces the source and target of the sequence, $\mathbf{T}\left(\overline{\mathcal{M}}\right)=\mathbf{S}\left(\mathcal{M}\right)$, $\mathbf{S}\left(\overline{\mathcal{M}}\right)=\mathbf{T}\left(\mathcal{M}\right)$, $T\left(\overline{A}\right)=S\left(A\right)$, and $S\left(\overline{A}\right)=T\left(A\right)$. Reversal is an involution,
$\overline{\overline{\mathcal{M}}}=\mathcal{M},$ (2.29)
that distributes over coarsening. It reverses the order of arguments when distributing over chaining (sometimes called “anti-involution” or “anti-isomorphism”),
$\overline{A\vee B}$ $=\overline{A}\vee\overline{B},$ (2.30)
$\overline{\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}}$ $=\overline{\mathcal{M}^{\mathrm{B}}}\cdot\overline{\mathcal{M}^{\mathrm{A}}}.$ (2.31)
$\overline{A\cdot B}$ $=\overline{B}\cdot\overline{A}.$ (2.32)
The reverse of a path is also called its adjoint.
Symmetric sequences are self-adjoint, i.e., they remain invariant under reversal. For symmetric $^{\circledcirc}\negthinspace\mathcal{M}\in\mathbb{\left[\mathbb{M}\right]}$ and $^{\circledcirc}\negthinspace A\in\mathcal{S}_{{}^{\circledcirc}\negthinspace\mathcal{M}}$ there is:
[…]
$\overline{{}^{\circledcirc}\negthinspace A}$ $\,=\,^{\circledcirc}\negthinspace A.$ (2.34)
2.3.6 Inverse operations unchaining and refinement
When we build a sequence of measurements individually, we have complete knowledge of all steps in its construction. Just as we can form new measurement sequences by chaining them together or coarsening the set of some detector outcomes, we can recover the prior sequences by simply not doing these operations. The same applies to paths once their underlying prior measurement setup has been recovered. In the model algebra, we reflect this by formally introducing inverse operations unchaining and refinement.
After chaining measurement sequences, $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}=\mathcal{M}^{\mathrm{C}}$ or result paths $A\cdot B=C$, we can start with $\mathcal{M}^{\mathrm{C}}$ or $C$ and formally recover the prior (noncommutative) arguments by left- and right-unchaining, $\setminus$ and //, respectively:
$/,\setminus$ $:\mathbb{\left[\mathbb{M}\right]}\times\mathbb{\left[\mathbb{M}\right]}\rightharpoonup\mathbb{\left[\mathbb{M}\right]},$
$\mathcal{M}^{\mathrm{C}}/\mathcal{M}^{\mathrm{B}}$ $=\left(\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}\right)/\mathcal{M}^{\mathrm{B}}=\mathcal{M}^{\mathrm{A}},$ (2.35)
$\mathcal{M}^{\mathrm{A}}\setminus\mathcal{M}^{\mathrm{C}}$ $=\mathcal{M}^{\mathrm{A}}\setminus\left(\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}\right)=\mathcal{M}^{\mathrm{B}},$ (2.36)
$/,\setminus$ $:\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\times\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\rightharpoonup\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}},$
$C/B$ $=\left(A\cdot B\right)/B=A,$ (2.37)
$A\setminus C$ $=A\setminus\left(A\cdot B\right)=B.$ (2.38)
Commuting or associating unchaining is generally undefined.
For coarsened paths $A\vee B=C$ (where $A,B\in\mathcal{S}_{\mathcal{M}}$, $C\in\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}}$), we start with $C$ and recover the prior argument by refinement $\wedge$:
$\wedge$ $:\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}}\times\mathcal{S}_{\mathcal{M}}\rightharpoonup\mathcal{S}_{\mathbb{\mathcal{M}}},$
$C\wedge B$ $=A,$ (2.39)
$C\wedge A$ $=B.$ (2.40)
Because coarsening does not depend on the order of its arguments, we do not need to distinguish between a left- and right-inverse operation. Just as for unchaining, commuting or associating refinement is generally undefined.
These inverses are partial binary operations only on a subset of allowable pairs of measurement sequences and result paths. We point out that inverses must nevertheless be partially well-defined, i.e., the operations are unique wherever they do exist [^5] A conceptual example would be squaring in positive whole numbers $\mathbb{Z}^{+}$, for which inverses are well-defined where they exist. In contrast, the inverse operation to squaring in the integers $\mathbb{Z}$ would not be well-defined in this sense: Where allowed, it generally permits two possible solutions (the only exception being 0).: We have full knowledge of the details of a given sequence and know which parts can be removed; and with that, we know what happens when removing some of these parts.
2.3.7 Insertion
Insertion is now introduced as an operation that expands a given sequence by adding measurements to a sequence or results to a path.
First, we define insertion in the most general case, adding a single measurement or result at a specified sequence index. Given a measurement sequence, $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]\in\left[\mathbb{M}\right]$, we insert a new measurement $\mathbf{M}_{\lambda}\in\mathbb{M}$ at position $j\in\left\{2\ldots\ell\right\}\in\mathbb{Z}^{+}$ by shifting the existing measurement $\mathbf{M}_{j}\in M$ at $j$ up by one index, and then adding the new measurement $\mathbf{M}_{\lambda}$ at the position occupied initially by $\mathbf{M}_{j}$:
$\downarrow$ $:\left[\mathbb{M}\right]\times\mathbb{M}\times\mathbb{Z}^{+}\rightarrow\left[\mathbb{M}\right],$
$\mathcal{M}\downarrow_{j}\mathbf{M}_{\lambda}$ $:=\left[\mathbf{M}_{1},\ldots,\mathbf{M}_{j-1},\mathbf{M}_{\lambda},\mathbf{M}_{j},\ldots,\mathbf{M}_{\ell}\right].$ (2.41)
The operation leaves the source and target of the original sequence $\mathcal{M}$ unchanged but increases its length by 1,
$\ell_{\mathcal{M}\downarrow_{j}\mathbf{M}^{\lambda}}=\ell_{\mathcal{M}}+1.$ (2.42)
Similarly, for measurement results $A=\left[A_{1},A_{2},\ldots,A_{\ell}\right]\in\mathcal{S}_{\mathcal{M}}$ we insert a new result $A_{\lambda}\in\mathbf{M}_{\lambda}$ into the path by:
$\downarrow$ $:\mathcal{S}_{\mathcal{M}}\rightarrow\mathcal{S}_{\mathcal{M}\downarrow_{j}\mathbf{M}_{\lambda}},$
$A\downarrow_{j}A_{\lambda}$ $:=\left[A_{1},\ldots,A_{j-1},A_{\lambda},A_{j},\ldots,A_{\ell}\right].$ (2.43)
For the sake of readability, this compact notation $\downarrow\,:\mathcal{S}_{\mathcal{M}}\rightarrow\mathcal{S}_{\mathcal{M}\downarrow_{j}\mathbf{M}_{\lambda}}$” accounts for inserting a new result $A_{\lambda}$ into the original path $\mathcal{S}_{\mathcal{M}}$ at position $j$, as well as a measurement $\mathbf{M}_{\lambda}$ that turns $\mathcal{M}$ into $\mathcal{M}\downarrow_{j}\mathbf{M}_{\lambda}$.
When inserting sequences of measurements and results, we use a similar notation as for the insertion of individual ones. Inserting a cyclic measurement sequence $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ after the sequence $\mathcal{M}^{\mathrm{J}}$ of a sequence chain $\mathcal{M}^{\mathrm{1}}\cdot\ldots\cdot\mathcal{M}^{\mathrm{L}}$ is written as:
$\downarrow$ $:\left[\mathbb{M}\right]\times\left[\mathbb{M}\right]\times\mathbb{Z}^{+}\rightarrow\left[\mathbb{M}\right],$
$\mathcal{M}\downarrow_{\mathrm{J}}{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ $=\mathcal{M}^{1}\cdot\mathcal{M}^{2}\cdot\ldots\cdot\mathcal{M}^{\mathrm{J}-1}\cdot{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\cdot\mathcal{M}^{\mathrm{J}}\cdot\ldots\cdot\mathcal{M}^{\mathrm{L}}.$ (2.44)
Chaining requires $\mathbf{T}\left(\mathcal{M}^{\mathrm{J}-1}\right)=\mathbf{S}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{T}\left({}^{\circ}\negthinspace\mathcal{M}^{\Lambda}\right)=\mathbf{S}\left(\mathcal{M}^{\mathrm{J}}\right)$ and this operation is only allowed for cyclic $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$. It increases sequence length by all elements of $^{\circ}\negthinspace\mathcal{M}^{\Lambda}$ that are not its source (or target):
$\ell_{\mathcal{M}\downarrow_{\mathrm{J}}{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}=\ell_{\mathcal{M}}+\ell_{{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}-1.$ (2.45)
Similarly for measurement results $A^{\Lambda}=\left[A_{1}^{\Lambda},A_{2}^{\Lambda},\ldots,A_{\ell_{\Lambda}}^{\Lambda}\right]\in\mathcal{S}_{{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}}$:
$\downarrow$ $:\mathcal{S}_{\mathcal{M}}\rightarrow\mathcal{S}_{\mathcal{M}\downarrow_{\mathrm{J}}{}^{\circ}\negthinspace\mathcal{M}^{\Lambda}},$
$A\downarrow_{\mathrm{J}}A^{\Lambda}$ $:=A^{1}\cdot\ldots A^{\mathrm{J}-1}\cdot A^{\Lambda}\cdot A^{\mathrm{J}}\ldots\cdot A^{\mathrm{L}}.$ (2.46)
3 Probability
So far, we have laid out an operational model with experiments that are sequences of measurements, possible detector outcomes, and paths that are sequences of measurement results when an experiment is performed. We have traced algebraic properties from select operations on these sequences: chaining, coarsening, their inverses, reversal, and insertions. In this model algebra, the coarsening operation is motivated from Quantum Mechanics: We want to permit detectors that fire on one of multiple possible, distinct underlying atomic outcome elements, where it does not need to be known which one exactly triggered the signal.
As a next step, we clarify the questions one may want to ask from the operational model: In Quantum Mechanics, the experimenter asks for probability distributions across possible measurement outcomes, given some preparation. The concept of probability is extraneous to our model; it is the experimenter who imposes it as mathematical construct in order to formulate the question. Nature, in turn, yields an answer given this constraint. If the experimenter is asking good questions, and the model represents nature, then the answers will be useful.
3.1 Paths have probability and establish closure
A general path $A=\left[A_{1},A_{2},\ldots,A_{\ell}\right]\in\mathcal{S}_{\mathcal{M}}$ over measurements $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]\in\left[\mathbb{M}\right]$, $\ell\geq 2$, will now be associated with a probability $P\in\left[0,1\right]\in\mathbb{R}$. Each measurement $\mathbf{M}_{j}$ has $\left|\mathbf{M}_{j}\right|$ detectors and a total of $\prod_{j=1}^{\ell}\left|\mathbf{M}_{j}\right|$ different paths can be observed over $\mathcal{S}_{\mathcal{M}}$. From repeated experiments using the same setup, a table can then be compiled that counts how often each allowable path occurred.
The probability of an individual path $A$ is defined as the conditional probability
$P$ $:\mathcal{S}_{\mathcal{M}}\rightarrow\left[0,1\right]\in\mathbb{R},$
$P\left(A\right)$ $:=\Pr\left(A_{\ell},A_{\ell-1},\ldots A_{2}\mid A_{1}\right).$ (3.1)
Conditionalization on the first result $S\left(A\right)=A_{1}$ ensures that the probability $P\left(A\right)$ is independent of the system’s history prior to the first measurement. We say that a path’s source establishes closure.
Note that we’re only making statements about the probability of transitions between measurement outcomes, not about the probabilities of outcomes by themselves. All measurement sequences are considered prepared by the first measurement result, which yields conditional probabilities for measurements thereafter.
3.2 Strong repeatability and factorization
Repeated measurements made from the identical detectors always produce the same result [^6] In this paper, we don’t consider interactions (or any other dynamics) to happen between measurements. Therefore, the wording “repeated,” “subsequent,” or “consecutive” for measurements always implies that nothing of consideration is happening in between.. In a sequence $\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,\mathbf{M}_{\ell}\right]$ that includes two consecutive identical measurements, $\mathbf{M}_{j}=\mathbf{M}_{j+1}$, the outcomes of $\mathbf{M}_{j}$ and $\mathbf{M}_{j+1}$ will always be identical, no matter the preparations at $\mathbf{M}_{1}$. With that, one of the $\mathbf{M}_{j},\mathbf{M}_{j+1}$ is redundant and can be removed from the calculation entirely. We call this strong repeatability.
Paths must start and end with atomic measurements and have atomic source and target outcomes. If a measurement sequence $\mathcal{M}^{\mathrm{A}}$ contains another atomic measurement, $^{\alpha}\mathbf{M}_{j}$ with $j\neq 1,\ell$, then we can decompose $\mathcal{M}^{\mathrm{A}}$ into two sequences, $\mathcal{M}^{\mathrm{B}}$ and $\mathcal{M}^{\mathrm{C}}$, effectively duplicating $^{\alpha}\mathbf{M}_{j}$ and designating it as $\mathbf{T}\left(B\right)$ and $\mathbf{S}\left(C\right)$, respectively:
$\mathcal{M}^{\mathrm{A}}$ $=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,{}^{\alpha}\mathbf{M}_{j},\ldots,\mathbf{M}_{\ell}\right],$
$\mathcal{M}^{\mathrm{B}}$ $:=\left[\mathbf{M}_{1},\mathbf{M}_{2},\ldots,{}^{\alpha}\mathbf{M}_{j}\right],$
$\mathcal{M}^{\mathrm{C}}$ $:=\left[{}^{\alpha}\mathbf{M}_{j},\mathbf{M}_{j+1},\ldots,\mathbf{M}_{\ell}\right],$
$\mathcal{M}^{\mathrm{A}}$ $=\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{C}}.$ (3.2)
This applies to paths accordingly:
$A$ $\in\mathcal{S}_{\mathcal{M}^{\mathrm{A}}},\,\,\,\,B\in\mathcal{S}_{\mathcal{M}^{\mathrm{B}}},\,\,\,\,C\in\mathcal{S}_{\mathcal{M}^{\mathrm{C}}},$
$A$ $=B\cdot C.$ (3.3)
With this, sequences $\mathcal{M}^{\mathrm{A}}$ and $A$ can always be unchained:
$\mathcal{M}^{\mathrm{B}}$ $=\mathcal{M}^{\mathrm{A}}/\mathcal{M}^{\mathrm{C}},$ (3.4)
$\mathcal{M}^{\mathrm{C}}$ $=\mathcal{M}^{\mathrm{B}}\setminus\mathcal{M}^{\mathrm{A}},$ (3.5)
$B$ $=A/C,$ (3.6)
$C$ $=B\setminus A.$ (3.7)
This process of unchaining can be repeated on $\mathcal{M}^{\mathrm{A}}$ until all measurements $\mathcal{M}^{\mathrm{J}}$, $\mathrm{J}=1\ldots\mathrm{L}$, and paths $A^{\mathrm{J}}$ are atomic only at their respective source and target:
$\mathcal{M}^{\mathrm{A}}$ $=\prod_{\mathrm{J}=1}^{\mathrm{L}}\mathcal{M}^{\mathrm{J}}=\mathcal{M}^{1}\cdot\mathcal{M}^{2}\cdot\ldots\cdot\mathcal{M}^{\mathrm{L}},$ (3.8)
$A$ $=\prod_{\mathrm{J}=1}^{\mathrm{L}}A^{\mathrm{J}}=A^{1}\cdot A^{2}\cdot\ldots\cdot A^{\mathrm{L}}.$ (3.9)
This process is called factorization. The individual $\mathcal{M}^{\mathrm{J}}$ and $\mathcal{A}^{\mathrm{J}}$ are called undecomposable (or ununchainable, if you wish).
3.3 Closure as Markov property: Histories don’t matter
Since the probability of a path is independent of the history of the system prior to its first measurement, factorization of paths corresponds to the Markov property for conditional probabilities in conventional probability theory,
$P\left(A\right)$ $=\Pr\left(A_{\ell_{\mathrm{L}}}^{\mathrm{L}},\ldots,A_{1}^{\mathrm{L}},A_{\ell_{\mathrm{L}-1}}^{\mathrm{L}-1},\ldots A_{1}^{\mathrm{L}-1},\ldots,A_{\ell_{2}}^{2},\ldots A_{1}^{2},A_{\ell_{1}}^{1},\ldots A_{2}^{1}\mid A_{1}^{1}\right).$
$=\prod_{\mathrm{J}=1}^{\mathrm{L}}\Pr\left(A_{\ell_{\mathrm{J}}}^{\mathrm{J}},A_{\ell_{\mathrm{J}}-1}^{\mathrm{J}},\ldots A_{2}^{\mathrm{J}}\mid A_{1}^{\mathrm{J}}\right)=\prod_{\mathrm{J}=1}^{\mathrm{L}}P\left(A^{\mathrm{J}}\right).$ (3.10)
We say that any atomic measurement establishes closure, such that histories do not matter at that point.
3.4 Weak repeatability
While strong repeatability requires identical measurements, we now also require weakly equivalent detectors to preserve the measurement of underlying atomic outcome elements, whether observed as an atomic result or assumed as part of a coarsened (non-atomic) result. This is called weak repeatability [^7] This term was introduced in [BCL1990] for a similar notion in the context of “generalized quantum mechanics”, which explored generalization of an observable compatible with operator quantum mechanics and quantum states..
Figure 3.1: In the example of an experiment $\mathcal{M}=\left[{}^{\alpha}\mathbf{M},{}^{\beta}\mathbf{M},{}^{\gamma}\mathbf{M},{}^{\alpha}\mathbf{M}\right],$ made from three weakly equivalent detectors $^{\alpha}\mathbf{M}\sim{}^{\beta}\mathbf{M}\sim{}^{\gamma}\mathbf{M}$, the probability of a sequence of measurement results is 11 if all detectors share the same underlying atomic outcome element (here: $m^{\prime}$ is part of all detectors of AA); otherwise, the probability is 0.
For example (figure 3.1), the measurements
$^{\alpha}\mathbf{M}$ $=\left\{\left\{m\right\},\left\{m^{\prime}\right\},\left\{m^{\prime\prime}\right\}\right\},$
$^{\beta}\mathbf{M}$ $=\left\{\left\{m,m^{\prime}\right\},\left\{m^{\prime\prime}\right\}\right\},$ (3.11)
$^{\gamma}\mathbf{M}$ $=\left\{\left\{m\right\},\left\{m^{\prime},m^{\prime\prime}\right\}\right\},$
are weakly equivalent, $^{\alpha}\mathbf{M}\sim{}^{\beta}\mathbf{M}\sim{}^{\gamma}\mathbf{M}$, since they share the same underlying atomic outcome elements $m,m^{\prime},m^{\prime\prime}$. Weak repeatability then preserves the overlapping measurement outcome $m^{\prime}$ in the path $A\in\mathcal{S_{M}}$:
$\mathcal{M}$ $=\left[{}^{\alpha}\mathbf{M},{}^{\beta}\mathbf{M},{}^{\gamma}\mathbf{M},{}^{\alpha}\mathbf{M}\right],$ (3.12)
$A$ $=\left[\left\{m^{\prime}\right\},\left\{m,m^{\prime}\right\},\left\{m^{\prime},m^{\prime\prime}\right\},\left\{m^{\prime}\right\}\right].$ (3.13)
Note that $A$ is the only path with nonzero probability over $\mathcal{M}$ that starts with $\left\{m^{\prime}\right\}$ since the underlying atomic outcome element $m^{\prime}$ is preserved in the experiment. All other paths starting with $\left\{m^{\prime}\right\}$ have probability zero and are hence impossible, e.g. $B,C\in\mathcal{S_{M}}$,
$B$ $=\left[\left\{m^{\prime}\right\},\left\{m^{\prime\prime}\right\},\left\{m^{\prime},m^{\prime\prime}\right\},\left\{m^{\prime}\right\}\right]\textrm{, or}$ (3.14)
$C$ $=\left[\left\{m^{\prime}\right\},\left\{m,m^{\prime}\right\},\left\{m\right\},\left\{m\right\}\right].$ (3.15)
Path $B$ is impossible because $m^{\prime}$ is not contained in the outcome $\left\{m^{\prime\prime}\right\}\in{}^{\beta}\mathbf{M}$, and path $C$ is impossible because $m^{\prime}$ is not contained in $\left\{m\right\}\in{}^{\gamma}\mathbf{M}$ (it does not matter that $m$ is in $\left\{m,m^{\prime}\right\}\in{}^{\beta}\mathbf{M}$). This gives us probabilities for these paths:
$P\left(A\right)$ $=\Pr\left(\left\{m^{\prime}\right\},\left\{m^{\prime},m^{\prime\prime}\right\},\left\{m,m^{\prime}\right\}\mid\left\{m^{\prime}\right\}\right)=1,$ (3.16)
$P\left(B\right)=P\left(C\right)$ $=0.$ (3.17)
Weak repeatability also allows inference given coarse-grained measurements: In an arbitrary sequence that somewhere contains $^{\beta}\mathbf{M}$ directly followed by $^{\alpha}\mathbf{M}$, once outcome $\left\{m,m^{\prime}\right\}\in{}^{\beta}\mathbf{M}$ is detected, the subsequent measurement $^{\alpha}\mathbf{M}$ will yield one of $\left\{m\right\}$ or $\left\{m^{\prime}\right\}$ with certainty (and never $\left\{m^{\prime\prime}\right\}$). Note that this inference applies only to consecutive weakly equivalent measurements.
3.5 Certain measurements
Measurements consisting of only a single detector will always measure the outcome from that detector. Such a detector is fully coarse-grained across all underlying atomic outcome elements, and the measurement result is certain.
Certain measurements [^8] The “no disturbance” postulate in [Goyal2014] used the term “trivial measurement” for the same concept. In this paper, we chose “certain” here to avoid confusion with our use of “trivial” for repeated identical measurements. may be inserted anywhere into any existing path without changing that path’s probability. Given any $\mathcal{M}\in\left[\mathbb{M}\right]$, $A\in\mathcal{S}_{\mathcal{M}}$, a fully coarse-grained measurement $^{\cup}\mathbf{M}\in\mathbb{M}$ with single outcome, $^{\cup}\mathbf{M}=\left\{{}^{\cup}M\right\}$, yields a certain measurement result, and we have for any $j\in\left\{2,\ldots,\ell_{\mathcal{M}}\right\}$:
$P\left(A\downarrow_{j}{}^{\cup}M\right)$ $=P\left(A\right).$ (3.18)
3.6 Impossible paths
Paths that contain different results from two subsequent identical measurements are impossible due to strong repeatability, i.e., they have a probability of 0. Likewise, paths with two subsequent weakly equivalent measurements are impossible if they do not share at least one underlying atomic outcome element in their respective results.
In both cases, the impossible paths contain at least one impossibility-generating pair (or short, IGP), a pair of consecutive weakly equivalent measurements, where the results do not contain a common possible atomic outcome element.
We now introduce impossible paths as a formal construct that is not realized in nature and derive specific properties in the model algebra:
  1. 1.
    Two impossible paths coarsened (or refined) always result in an impossible path,
  2. 2.
    coarsening of a possible path with an impossible path always yields a possible path,
  3. 3.
    refinement of a possible path by an impossible path always yields a possible path.
Note that impossible paths are formally valid in the model algebra and operations thereon; they merely have probability zero. This is distinct from operands in disallowed operations, which cannot be executed in the partial model algebra in principle.
First, to prove that two impossible paths coarsened (or refined) must always yield an impossible path (“1.”):
Proof.
Coarsening may only modify exactly one element in a measurement sequence. There are two cases: (1) the sequence step modified under coarsening is not part of an IGP, or (2) the sequence step is one of an IGP. In case (1), the resulting path must also be impossible, since no IGP is modified. In the case of (2), only one of the two measurements in the IGP is modified. If the result would make the IGP possible, this would mean that the resulting path contains the same underlying atomic outcome element in both measurement results of the original IGP. This would be a contradiction because only one of the results is modified, and both initial paths are impossible and therefore do not contain any shared underlying atomic outcome elements at the IGP. A similar argument applies to refinement. ∎
Now proving that coarsening a possible path with an impossible path always yields a possible path (“2.”):
Proof.
Coarsening two paths always results in a path that has a single sequence step modified. There, the coarsened result is built over the initial paths’ underlying atomic outcome elements. Because this always includes all outcome elements from the initial paths, it will contain all elements from the possible path. ∎
Finally, the proof that refining a possible path by an impossible path always yields another possible path (“3.”):
Proof.
Refinement of a possible path by an impossible path must always happen at the single IGP of the impossible path since, otherwise, both paths would contain that IGP and be impossible. At the IGP, the possible and impossible paths have one measurement result in common and differ by the other. At those two steps, the possible path must share at least one underlying atomic outcome element across both measurement results. If the impossible path were to remove these shared outcome elements from the possible path, the impossible path would have also had to contain these outcome elements in the IGP. This would make the impossible path possible and is a contradiction. ∎
3.7 Equivalent paths and redundancy
Per strong repeatability, one of two consecutive equal measurements may be removed from the calculation. Duplicating measurements could, therefore, be viewed as some identity operation under chaining in the model algebra. Similarly, elements from the space of impossible paths act on other paths as identity elements under coarsening.
We now formally define a class of equivalent paths from these two scenarios, such that any two equivalent paths can be exchanged at will, and treated as equal in the calculations to come.
Two paths are equivalent if they can be changed into one another …
  1. 1.
    …through coarsening or refinement with impossible paths; or
  2. 2.
    … through chaining, unchaining, insertion, or removal of identical measurement results from consecutive equal measurements.
Given all possible paths that are equivalent under this class, there is a unique minimal possible path, now called nonredundant, such that every other path equivalent to it contains some redundancy.
Proof.
Per (1), coarsening and refinement are generally only allowed if the operands differ by one measurement result at a single sequence step. At all other sequence steps, the paths must be identical. Coarsening a possible path with an impossible path, therefore, enlarges the possible path by outcome elements of an IGP. Repeated refinement, in turn, reduces the number of elements from a path. Once all elements from any IGP are removed from a possible path, such a path is uniquely minimal under refinement. Per (2), identical measurement results can be trivially chained, unchained, inserted, or removed. A path is uniquely minimal under removal once it contains no more consecutive duplicates. Taken together, unchaining from (1) and removal from (2) can be repeated on any given possible path until it is uniquely minimal in both cases. The resulting path is, therefore, possible, unique under the class, and contains no redundancy. ∎
We extend the definition of symmetric (palindromic) paths accordingly to paths that are symmetric in their nonredundant form.
For notation, two equivalent paths $A,B$ may be written $A\sim B$ when emphasizing the possibility of different redundancy in $A$ and $B$; but generally, $A=B$ suffices since equivalent paths are effectively equal.
3.8 Extending coarsening for equivalent paths
When we first introduced coarsening in section 2.3.1, we defined $\mathcal{M}^{\vee\left\{j\right\}}$ as the sequence that differs from a given sequence $\mathcal{M}$ by replacing exactly one of its measurements, $\mathbf{M}_{j}$, with a coarsened version $^{\vee}\mathbf{M}_{j}$. The coarsening operation $A\vee B=C$ is then allowable only if $A,B\in\mathcal{S}_{\mathcal{M}}$ and $C\in\mathcal{S}_{\mathcal{M}^{\lor\left\{j\right\}}}$, and the underlying atomic outcome elements of $A$ and $B$ at $j$ do not overlap.
Most general paths do not meet these conditions and cannot be coarsened. However, even if these conditions are met for specific paths $A,B,C$, they are generally not satisfied for paths equivalent to these under the redundancy class. Therefore, we extend the partial operation coarsening $A\vee B=C$ to be allowable if the operation’s prerequisites can be met for at least one equivalent member representing each of the operands. This is consistent with treating equivalent paths as equal while evaluating expressions in the model algebra on suitable members representing that equivalence class.
3.9 Connection to classical probability
Insertion of a certain measurement does not change a path’s probability, whereas insertion of a measurement that generates a new IGP forces the path’s probability to 0. These operations are conceptually similar to conventional probability, where insertion of a certain or impossible event into an existing set of independent events has the same effect on the joint probability for these events [^9] Note that insertion of a certain measurement into a path does not have to generate redundancy and is therefore not necessarily trivial, as it would be in conventional probability..
Furthermore, in an experiment with two measurements, $\mathcal{M}^{\mathbf{D}}=\left[\mathbf{M}_{1},\mathbf{M}_{2}\right]\in\left[\mathbb{M}\right]$, the sum of probabilities over all possible paths $D_{r}$ that start with $M_{1}\in\mathbf{M}_{1}$ and end on one of the $M_{2}^{r}\in\mathbf{M}_{2}$ must add up to $1$, since detectors are distinct and non-filtering:
$D_{r}$ $:=\left[M_{1},M_{2}^{r}\right]\in\mathcal{S}_{\mathcal{M^{\mathbf{D}}}},$ (3.19)
$\sum_{\forall r}P\left(D_{r}\right)$ $=\sum_{\forall r}\Pr\left(M_{2}^{r}\mid M_{1}\right)=1.$ (3.20)
4 Amplitude algebra
Equipped with an axiomatic set of partial operations from an elementary model and specific questions we want to ask from that model, we can now choose the concept of amplitudes to represent answers to our questions. We will support this choice with a simplicity argument, show that it is valid without loss of generality for the task at hand, and not in conflict with the model in general. This executes our stated, programmatic goal from the introduction: to avoid premature optimization of the axiomatic model algebra, for which we may not have enough structure to attempt classification; to bring forward a choice that allows answering specific questions; and to provide the basis for complete classification of consequences from that choice, including an explanation towards the utility of the complex numbers in canonical Quantum Mechanics.
4.1 Representation space
An amplitude $\mathbf{a}$ is an element in an axiomatic space $\mathbb{A}$ that is now constructed to quantify probability behavior under operations of the model algebra. Given an experiment $\mathcal{M}\in\left[\mathbb{M}\right]$ and path $A\in\mathcal{S}_{\mathcal{M}}$, a function $\phi$ represents that path $A$ as an amplitude $\mathbf{a}\in\mathbb{A}$:
$\phi$ $:\mathcal{S}_{\left[\mathbb{M}\right]}\rightarrow\mathbb{A},$
$\mathcal{M}$ $\in\left[\mathbb{M}\right],\,\,\,\,A\in\mathcal{S}_{\mathcal{M}},\,\,\,\,\mathbf{a}\in\mathbb{A},$ (4.1)
$\mathbf{a}$ $=\phi\left(A\right).$
The representation function $\phi$ is a well-defined map of outcomes of an experiment $\mathcal{S}_{\left[\mathbb{M}\right]}$ into the space of amplitudes $\mathbb{A}$. Equivalent paths $A,B\in\mathcal{S}_{\mathcal{M}}$ have identical representation,
$A\sim B\quad\rightarrow\quad\phi\left(A\right)=\phi\left(B\right).$ (4.2)
The representation map $\phi$ does not have to be invertible, and not every value in $\mathbb{A}$ has to represent an allowable path in $\mathcal{S}_{\left[\mathbb{M}\right]}$.
The probability $P$ of a path is identical to the $p$-function of the amplitude representing that path:
$p$ $:\mathbb{A}\rightarrow\mathbb{R},$
$p\left(\phi\left(A\right)\right)$ $=P\left(A\right).$ (4.3)
For amplitudes that do not represent allowable paths, the image of the $p$-function is not bounded to the probability interval $\left[0,1\right]\in\mathbb{R}$ but formally extended to all of $\mathbb{R}$.
The concept of disallowed subspaces in amplitude representation of nature is nothing new: In canonical Quantum Mechanics, distinct states of some system may be represented as amplitudes in the space of complex numbers, $\mathbb{C}$. There, the Born rule allows to compute the probability of finding the system in one of these states, by forming the squared norm of their respective amplitudes. Because probabilities greater than $1$ are nonsensical, any number in $\mathbb{C}$ with a norm greater than $1$ can never represent an allowable state in these systems. However, all complex numbers are formally useful in the calculation, if abstract.
By tracing algebraic operations on amplitudes in $\mathbb{A}$ from a model algebra, regardless of whether such operations would make sense in the original model, we make a subtle but important choice:
  • Any valid operation in the model algebra must have a well-defined algebraic representation on amplitudes, and
  • any permissible operation on amplitudes must not conflict with allowable operations in the model.
These two constraints allow the freedom of valid algebraic properties on amplitudes that cannot be traced from the model in principle. All valid properties on amplitudes – traceable or not – must have utility when answering questions the observer brings to the experiment, namely, asking for probabilities.
4.2 Core algebraic properties
The model algebra of paths $\mathcal{S}_{\left[\mathbb{M}\right]}$ over experiments $\left[\mathbb{M}\right]$ has coarsening $\vee:\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\times\mathcal{S}_{\mathbb{\left[\mathbb{M}\right]}}\rightharpoonup\mathcal{S}_{\left[\mathbb{M}\right]}$ and chaining $\cdot:\mathcal{S}_{\left[\mathbb{M}\right]}\times\mathcal{S}_{\left[\mathbb{M}\right]}\rightharpoonup\mathcal{S}_{\left[\mathbb{M}\right]}$ as its elementary partial operations, abbreviated $\left\langle\mathcal{S}_{\left[\mathbb{M}\right]},\vee,\cdot\right\rangle$. This is now paired with algebraic operations $\oplus$ and $\odot$ on the amplitude space, $\left\langle\mathbb{A},\oplus,\odot\right\rangle$ using some mapping $\phi$:
$\left\langle\mathcal{S}_{\left[\mathbb{M}\right]},\vee,\cdot\right\rangle\overset{\phi}{\longmapsto}\left\langle\mathbb{A},\oplus,\odot\right\rangle.$ (4.4)
In contrast to the model algebra, binary operations in the amplitude algebra are defined on all elements of the space. The symbols $\oplus$ and $\odot$ are chosen suggestively, termed amplitude addition and multiplication, respectively. At this point, however, these are just placeholder symbols and terms.
For some experiment $\mathcal{M}\in\left[\mathbb{M}\right]$ and paths $A,B\in\mathcal{S}_{\mathcal{M}}$ that are allowed to be coarsened, the amplitude mapping $\phi:\mathcal{S}_{\left[\mathbb{M}\right]}\rightarrow\mathbb{A}$ has the property:
$\phi\left(A\vee B\right)=\phi\left(A\right)\oplus\phi\left(B\right).$ (4.5)
Coarsening is commutative and partially associative, from which we trace for any $\mathbf{a},\mathbf{b},\mathbf{c}\in\mathbb{A}$:
$\oplus$ $:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{A},$
$\left(\mathbf{a}\oplus\mathbf{b}\right)\oplus\mathbf{c}$ $=\mathbf{a}\oplus\left(\mathbf{b}\oplus\mathbf{c}\right),$ (4.6)
$\mathbf{a}\oplus\mathbf{b}$ $=\mathbf{b}\oplus\mathbf{a}.$ (4.7)
For chaining, we have accordingly:
$\phi\left(A\cdot B\right)=\phi\left(A\right)\odot\phi\left(B\right).$ (4.8)
Closure as Markov property for probabilities (equation 3.10) extends to amplitudes:
$P\left(A\cdot B\right)$ $=P\left(A\right)P\left(B\right)$
$=p\left(\phi\left(A\cdot B\right)\right)$ $=p\left(\phi\left(A\right)\odot\phi\left(B\right)\right)$ (4.9)
$=p\left(\mathbf{a}\odot\mathbf{b}\right)$ $=p\left(\mathbf{a}\right)p\left(\mathbf{b}\right).$
Chaining is not required to be commutative but associative. This becomes:
$\odot$ $:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{A},$
$\left(\mathbf{a}\odot\mathbf{b}\right)\odot\mathbf{c}$ $=\mathbf{a}\odot\left(\mathbf{b}\odot\mathbf{c}\right).$ (4.10)
Chaining distributes over coarsening, which holds in amplitude representation as well:
$\left(\mathbf{a}\oplus\mathbf{b}\right)\odot\mathbf{c}$ $=\mathbf{a}\odot\mathbf{c}\oplus\mathbf{b}\odot\mathbf{c},$ (4.11)
$\mathbf{a}\odot\left(\mathbf{b}\oplus\mathbf{c}\right)$ $=\mathbf{a}\odot\mathbf{b}\oplus\mathbf{a}\odot\mathbf{c}.$ (4.12)
These are the core algebraic relations (S1) through (S5) from [GKS2010].
4.3 Involution
[…]
$\overline{\overline{\mathbf{a}}}$ $=\mathbf{a},$ (4.13)
$\overline{\mathbf{a}\oplus\mathbf{b}}=\overline{\mathbf{b}\oplus\mathbf{a}}$ $=\overline{\mathbf{a}}\oplus\overline{\mathbf{b}}=\overline{\mathbf{b}}\oplus\overline{\mathbf{a}},$ (4.14)
$\overline{\mathbf{a}\odot\mathbf{b}}$ $=\overline{\mathbf{b}}\odot\overline{\mathbf{a}}.$ (4.15)
Self-adjoint paths $^{\circledcirc}\negthinspace A$ are represented by amplitudes $\mathbf{d}=\phi\left(\,{}^{\circledcirc}\negthinspace A\right)$ that are invariant under involution, $\overline{\mathbf{d}}=\mathbf{d}$.
4.4 Impossible paths as additive identity
Section 3.6 introduced impossible paths as a formal construct for sequences of measurement results with probability 0. They contain at least one pair of results from weakly equivalent measurements that do not share an underlying atomic outcome element. We observe that impossible paths as a whole have properties similar to an additive identity under coarsening: Coarsening two impossible paths always results in another impossible path, coarsening a possible path with an impossible path always yields a possible path, and coarsening two possible paths always yields a possible path.
We, therefore, now choose to represent all impossible paths by the same amplitude, $\mathbf{0}\in\mathbb{A}$, and to be the unique additive identity on amplitudes. Given an arbitrary possible path $A$ with $\mathbf{a}=\phi\left(A\right)$, there is:
$\mathbf{0}\oplus\mathbf{0}$ $=\mathbf{0},$ (4.16)
$\mathbf{0}\oplus\mathbf{a}=\mathbf{a}\oplus\mathbf{0}$ $=\mathbf{a}.$ (4.17)
The probability for any impossible path is 0, which applies to their amplitude representation as well:
$p\left(\mathbf{0}\right)$ $=0.$ (4.18)
Paths that are possible, in contrast, have nonzero probability. For allowable paths path $B$ with $\mathbf{b}=\phi\left(B\right)$, we can therefore infer
$p\left(\phi\left(B\right)\right)$ $=0\quad\rightarrow\quad\mathbf{b}=\mathbf{0}.$ (4.19)
Note that we cannot make assertions about elements in $\mathbb{A}$ that do not represent allowable paths. If there are elements $\mathbf{c}\in\mathbb{A}$ where $p\left(\mathbf{c}\right)=0$ but $\mathbf{c}\neq\mathbf{0}$, then such $\mathbf{c}$ must never represent allowable paths in the model algebra.
Coarsening with an impossible path is a redundancy operation between equivalent paths per section 3.7. This is consistent with the algebraic constraint imposed on amplitudes here: Equivalent paths are exchangeable in the calculation at will and can therefore be represented by the same amplitude, $\mathbf{0}$.
4.5 Trivial paths as multiplicative identity
A trivial path is a sequence of results from an experiment that consists only of equal measurements. Per strong repeatability, all measurement results must be identical. Like coarsening with an impossible path, chaining with a trivial path is a redundancy operation.
We formally choose the model operation of chaining with a trivial path to be represented by an amplitude, $\mathbf{1}\in\mathbb{A}$, such that for any path $A$ with $\mathbf{a}=\phi\left(A\right)$, there is:
$\mathbf{a}\odot\mathbf{1}=\mathbf{1}\odot\mathbf{a}$ $=\mathbf{a}.$ (4.20)
Because $\mathbf{1}$ is a left- and right-identity, and amplitude multiplication $\odot$ is associative, $\mathbf{1}$ is a multiplicative identity for all $\mathbf{a}\in\mathbb{A}$, and
$p\left(\mathbf{1}\right)=1.$ (4.21)
In the model algebra, a trivial path $^{\mathbf{1}}\negthinspace A$ can be chained to any of the $A^{\mathrm{J}}$ of a path factorization $A=\prod_{\mathrm{J}=1}^{\mathrm{L}}A^{\mathrm{J}}=A^{1}\cdot A^{2}\cdot\ldots\cdot A^{\mathrm{L}}$, as long as all measurements in $^{\mathbf{1}}\negthinspace A$ are the same as the target of $A^{\mathrm{J}}$ (which is also the source of $A^{\mathrm{J}+1}$, of course):
$A^{\prime}$ $=A^{1}\cdot\ldots A^{\mathrm{J}}\cdot\,^{\mathbf{1}}\negthinspace A\cdot A^{\mathrm{J}+1}\ldots\cdot A^{\mathrm{L}},$ (4.22)
$\phi\left(A^{\prime}\right)$ $=\phi\left(A^{1}\right)\odot\ldots\odot\phi\left(A^{\mathrm{J}}\right)\odot\mathbf{1}\odot\phi\left(A^{\mathrm{J}+1}\right)\odot\ldots\odot\phi\left(A^{\mathrm{L}}\right)=\phi\left(A\right).$ (4.23)
This correspond to insertion of $^{\mathbf{1}}\negthinspace A$,
$A^{\prime}=A\downarrow_{\mathrm{J}}\,^{\mathbf{1}}\negthinspace A.$ (4.24)
Because chaining with a trivial path is a redundancy operation, the paths $A^{\prime}$ and $A$ are equal, and the length of their nonredundant form is not changed. This is consistent with representation by a multiplicative unit $\mathbf{1}=\phi\left(\,{}^{\mathbf{1}}\negthinspace A\right)$. It also means that the insertion index $\mathrm{J}$ in the expression $A\downarrow_{\mathrm{J}}\,^{\mathbf{1}}\negthinspace A$ is superfluous: Wherever it is allowable to insert $^{\mathbf{1}}\negthinspace A$, it results in an equivalent path. We therefore write:
$A\sim A\downarrow\,^{\mathbf{1}}\negthinspace A.$ (4.25)
In contrast, note that insertions of a single, certain measurement $A\downarrow_{j}{}^{\cup}M$, as in section 3.5, do not correspond to chaining in the model algebra and, therefore, have no algebraic representation. Even though this operation does not change the probability of the path, $P\left(A\downarrow_{j}{}^{\cup}M\right)=P\left(A\right)$, its amplitude $\phi\left(A\right)$ may be different from $\phi\left(A\downarrow_{j}{}^{\cup}M\right)$.
4.6 Invertibility and algebraic closure
What led to a coarsened path can be refined, and what led to a chained path can be unchained. We argued for partially well-defined inverse operations in the model algebra in section 2.3.6, where we have complete information on all detectors, their possible measurement outcomes, and actual measurements. In this section, we now show that additive invertibility must always hold in amplitude representation and that multiplicative invertibility must hold for paths with nonzero probability. We postulate algebraic closure for the formal existence of inverse elements in $\mathbb{A}$ that represent these inverse operations.
4.6.1 Additive inverse
For every nonzero element $\mathbf{a}\in\mathbb{A}$, $\mathbf{a}\neq\mathbf{0}$ that represents possible path $A\in\mathcal{S}_{\mathcal{M}}$, there must be an operation $\ominus$ that is inverse to $\oplus$ such that
$\left(\mathbf{0}\oplus\mathbf{a}\right)\ominus\mathbf{a}=\mathbf{0}.$ (4.26)
This follows directly from coarsening and subsequent refinement in the model algebra (where $\left(A\vee B\right)\wedge B=A$ per equation 2.39) and from $p\left(\left(\mathbf{0}\oplus\mathbf{a}\right)\ominus\mathbf{a}\right)=0$.
Next to the existence of an inverse operation to addition, algebraic closure requires the formal existence of inverse elements $\mathbf{a}\in\mathbb{A}$ representing this operation, such that $\mathbf{a}\oplus\left(-\mathbf{a}\right)=\mathbf{0}$ and therefore $p\left(\mathbf{a}\oplus\left(-\mathbf{a}\right)\right)=0$. Note that neither $\mathbf{a}$ nor the expression $\mathbf{a}\oplus\left(-\mathbf{a}\right)$ represent paths or path operations from the model algebra anymore; instead, these expressions are merely formally valid in amplitude representation in a way that abstracts some earlier expressions that did – at first – trace from path operations.
For testing uniqueness, we assume there exists a different amplitude $\mathbf{a}^{\prime}\in\mathbb{A}$ such that $\mathbf{a}\neq\mathbf{a}^{\prime}$ but $\left(\mathbf{0}\oplus\mathbf{a}\right)\ominus\mathbf{a}^{\prime}=\mathbf{0}$ as well. However, since $\mathbf{0}\oplus\mathbf{a}$ represents a possible path, refinement by another path can only yield an impossible path if both are equivalent up to redundancy and therefore have the same amplitude. Since that is not the case, $\left(\mathbf{0}\oplus\mathbf{a}\right)\ominus\mathbf{a}^{\prime}$ can never be zero for $\mathbf{a}\neq\mathbf{0}$ and $\mathbf{a}\neq\mathbf{a}^{\prime}$. Because amplitude representation must not create expressions that conflict with the model, such assumed $\mathbf{a}^{\prime}$ could not exist, and $\ominus\mathbf{a}$ is unique for any given nonzero $\mathbf{a}$.
The edge case $\mathbf{a}=\mathbf{a}^{\prime}=\mathbf{0}$ has $\left(\mathbf{0}\oplus\mathbf{0}\right)\ominus\mathbf{0}=\mathbf{0}$, which is always true. Coarsening and refinement between impossible paths always yield another impossible path. The unique additive inverse for $\mathbf{a}=\mathbf{0}$, therefore, is $\mathbf{0}$ itself.
Existence, uniqueness, and algebraic closure together make amplitude addition invertible in $\mathbb{A}$ for any $\mathbf{a}\in\mathbb{A}$.
4.6.2 Multiplicative inverse
[…]
$\left(\mathbf{1}\odot\mathbf{a}\right)\oslash\mathbf{a}$ $=\mathbf{1},$ (4.27)
$\mathbf{a}\obslash\left(\mathbf{a}\odot\mathbf{1}\right)$ $=\mathbf{1}.$ (4.28)
This follows from chaining and subsequent left- and right-unchaining in the model algebra (where $\left(A\cdot B\right)/B=A$ per equation 2.37, and $A\setminus\left(A\cdot B\right)=B$ per equation 2.38).
Algebraic closure then requires the formal existence of inverse elements $^{-1}\mathbf{a},\mathbf{a}^{-1}\in\mathbb{A}$ such that
$\left(\mathbf{1}\odot\mathbf{a}\right)\odot\mathbf{a}^{-1}$ $=\mathbf{1},$ (4.29)
$^{-1}\mathbf{a}\odot\left(\mathbf{a}\odot\mathbf{1}\right)$ $=\mathbf{1}.$ (4.30)
Since $\mathbf{1}\odot\mathbf{a}=\mathbf{a}\odot\mathbf{1}$ and $\odot$ is associative, these left- and right-inverse operations must be the same and can formally be represented by the same element under algebraic closure,
$^{-1}\mathbf{a}=\mathbf{a}^{-1}.$ (4.31)
Closure in the model algebra yields the Markov property for probability (equation 3.10), which for the $p$-function formally extends to the closure space of $\mathbb{A}$ under amplitude multiplication:
$p\left(\mathbf{a}\odot\mathbf{a}^{-1}\right)=p\left(\mathbf{a}\right)p\left(\mathbf{a}^{-1}\right)$ $=1.$ (4.32)
Just as the other expressions involving inverse elements, the $p$-function of a multiplicative inverse is merely a formal extension of $p\left(\mathbf{a}\right)$ to elements of the closure space. The product composition property $p\left(\mathbf{a}\odot\mathbf{b}\right)=p\left(\mathbf{a}\right)p\left(\mathbf{b}\right)$ is formally carried forward for inverses, $\mathbf{b}\equiv\mathbf{a}^{-1}$. With this, for $p\left(\mathbf{a}\right)p\left(\mathbf{a}^{-1}\right)=1$ to hold and for inverses $\mathbf{a}^{-1}$ to exist, $p\left(\mathbf{a}\right)$ and $p\left(\mathbf{a}^{-1}\right)$ must be nonzero.
In order to argue for uniqueness, we assume some nonzero $\mathbf{a}^{\prime}\in\mathbb{A}$, $\mathbf{a}^{\prime}\neq\mathbf{0}$ such that $\mathbf{a}\neq\mathbf{a}^{\prime}$ but $\left(\mathbf{1}\odot\mathbf{a}\right)\oslash\mathbf{a}^{\prime}=\mathbf{1}$. Just as in the additive case, $\mathbf{a}^{\prime}\neq\mathbf{a}$ must correspond to a path $A^{\prime}$ that differs from $A$ in its nonredundant form. Given a path $^{\mathbf{1}}B$ with unit amplitude $\phi\left({}^{\mathbf{1}}B\right)=\mathbf{1}$, unchaining $\left({}^{\mathbf{1}}B\cdot A\right)/A^{\prime}$ requires the factorization of $A$ to end on the factorization of $\overline{A^{\prime}}$, up to equivalence. For $\mathbf{a}^{\prime}\neq\mathbf{a}$ to hold, $A$ must, therefore, be longer than $A^{\prime}$, with a remainder $\delta A=A/A^{\prime}$ and $\phi\left(\delta A\right)\neq\mathbf{1}$. However, since $\phi\left(\delta A\right)=\phi\left(A/A^{\prime}\right)=\mathbf{a}\oslash\mathbf{a}^{\prime}\neq\mathbf{1}$, this contradicts the requirement $\left(\mathbf{1}\odot\mathbf{a}\right)\oslash\mathbf{a}^{\prime}=\mathbf{1}$, and this case is disallowed. The same argument applies to left-unchaining. Just as for addition, expressions in amplitude representation must never create valid expressions that conflict with the model. This makes multiplicative inverses of amplitudes representing allowable paths unique as well.
Therefore, we established the existence, uniqueness, and algebraic closure of multiplicative inverses of amplitudes $\mathbf{a}\in\mathbb{A}$, as long as $p\left(\mathbf{a}\right)\neq 0$.
4.7 Scalars
We now postulate the existence of a real-valued subspace of the amplitude algebra that is spanned by the multiplicative identity amplitude $\mathbf{1}$,
$\mathbb{R}\cong\mathbb{R}\mathbf{1}\subset\mathbb{A}.$ (4.33)
Elements $\lambda\in\mathbb{R}$ are called scalars and exist formally separate from amplitudes. We support this choice from conceptual simplicity: The probabilities we are ultimately looking for are quantified using real numbers between 0 and $1$. In this paper, the mathematical mechanism to obtain probabilities uses the more faceted construct of amplitudes. By embedding the reals into these amplitudes by postulate, we aim to keep amplitudes somewhat similar to probabilities, while exposing additional facets as algebraic variations thereof.
Subsequently, we postulate for amplitudes $\lambda\mathbf{1}\in\mathbb{R}\mathbf{1}$ to represent symmetric (palindromic) paths in the model. This choice is supported by internal consistency, where we will show that resulting properties in the amplitude algebra either trace from the model or do not conflict with it.
Whether these postulates are good choices cannot be answered mathematically. Instead, their utility has to be shown in the experiment. You could pointedly say that here we are doing physics.
Two symmetric paths $^{\circledcirc}\negthinspace B,\,^{\circledcirc}C\in\mathcal{S}_{{}^{\circledcirc}\negthinspace\mathcal{M}}$ with amplitudes $\mathbf{b}=\phi\left(\,{}^{\circledcirc}\negthinspace B\right),\mathbf{c}=\phi\left(\,{}^{\circledcirc}C\right)$,
$\overline{\,{}^{\circledcirc}\negthinspace B}$ $=\,^{\circledcirc}\negthinspace B,$ $\overline{\,{}^{\circledcirc}C}$ $=\,^{\circledcirc}C,$ (4.34)
$\overline{\mathbf{b}}$ $=\mathbf{b},$ $\overline{\mathbf{c}}$ $=\mathbf{c},$ (4.35)
correspond to real scalars $\lambda,\mu\in\mathbb{R}$:
$\mathbf{b}$ $:=\lambda\mathbf{1}\equiv\lambda,$ $\mathbf{c}$ $:=\mu\mathbf{1}\equiv\mu.$ (4.36)
Since scalars formally exist as separate entities from amplitudes, they may be notated separately and pulled up front by convention.
Chaining their paths then follows the same rules as for the remaining multiplicative identity. For example, given some general path $A$ with $\mathbf{a}=\phi\left(A\right)$, chaining with paths representing scalars is an insertion wherever allowable:
$\phi\left(A\downarrow\,^{\circledcirc}\negthinspace B\right)$ $=\lambda\,\phi\left(A\right),$ (4.37)
$\phi\left(A\downarrow\,^{\circledcirc}C\right)$ $=\mu\,\phi\left(A\right),$
$\mathbf{a}\odot\mathbf{b}=\mathbf{a}\odot\lambda\mathbf{1}$ $=\lambda\mathbf{a}=\lambda\mathbf{1}\odot\mathbf{a}=\mathbf{b}\odot\mathbf{a},$ (4.38)
$\mathbf{a}\odot\mathbf{b}\odot\mathbf{c}$ $=\lambda\mu\mathbf{a}=\mathbf{b}\odot\mathbf{a}\odot\mathbf{c}=\ldots\textrm{(all permutations)}.$ (4.39)
The zero scalar, $\nu=0$, is the annihilator under multiplication that forces the probability of any sequence to 0 and must therefore effect a zero amplitude when inserted:
$0\mathbf{a}$ $=\mathbf{0},$ (4.40)
[…]
With commutativity given by postulate of the real numbers, we have:
$\overline{\lambda\mathbf{1}\odot\mu\mathbf{1}}=\overline{\mu\mathbf{1}}\odot\overline{\lambda\mathbf{1}}=\mu\mathbf{1}\odot\lambda\mathbf{1}=\lambda\mathbf{1}\odot\mu\mathbf{1}.$ (4.42)
That is, the product of two amplitudes representing a symmetric path, $\sigma:=\lambda\mu\equiv\lambda\mathbf{1}\odot\mu\mathbf{1}$, is also self-adjoint, making the algebraic subspace closed under chaining. This identifies self-adjoint amplitudes with the reals,
$\left\{\sigma\mid\sigma\in\mathbb{R}\right\}\cong\left\{\mathbf{c}\mid\mathbf{c}\in\mathbb{A},\,\overline{\mathbf{c}}=\mathbf{c}\right\},$ (4.43)
independently of whether the represented path in the model is symmetric or not.
Because involution is linear,
$\overline{\lambda\mathbf{1}\odot\mathbf{a}}=\lambda\mathbf{1}\odot\overline{\mathbf{a}}\equiv\lambda\overline{\mathbf{a}}=\overline{\lambda\mathbf{a}},$ (4.44)
it is also continuous [^10] Note that Cauchy’s functional equation $\overline{\mathbf{a}\oplus\mathbf{b}}=\overline{\mathbf{a}}\oplus\overline{\mathbf{b}}$ for a continuous involution would allow us to deduce linearity. Here, we can start by knowing the stronger property, linearity, which always infers continuity over the reals..
Imposing a relatively strong choice, as is done here, risks constraining allowable amplitude algebras to cases that no longer exhibit interesting or useful structure. Given all possible permutations over a finite set of arbitrary measurements and detectors, the percentage of symmetric permutations generally decreases as the number of different detectors grows. Correspondingly, the subspace of the amplitude algebra on which we impose the real-number representation postulate becomes smaller the more variety is allowed. Some follow-on constraints derive from this, for example, for chains of symmetric paths to be represented by the reals as well necessarily. However, these cases remain limited in the overall scope of infinite variety and, therefore, for this postulate to not overconstrain model representation.
4.7.1 Continuous $p$-function
With closure of the $p$-function,
$p\left(\lambda\mathbf{a}\right)=p\left(\lambda\mathbf{1}\odot\mathbf{a}\right)=p\left(\lambda\mathbf{1}\right)p\left(\mathbf{a}\right),$ (4.45)
there exist functions $f:\mathbb{R}\rightarrow\mathbb{R}$ which only depend on $\lambda,\mu\in\mathbb{R}$:
$f\left(\lambda\right)$ $:=p\left(\lambda\mathbf{1}\right)=\frac{p\left(\lambda\mathbf{1}\odot\mathbf{a}\right)}{p\left(\mathbf{a}\right)},$ (4.46)
$f\left(\lambda\mu\right)$ $=f\left(\lambda\right)f\left(\mu\right).$ (4.47)
For continuous $f$, the general nonconstant solutions of this functional equation [^11] The constant solutions $f_{0}\left(\lambda\right)=0$ and $f_{1}\left(\lambda\right)=1$ do not give any information on probabilities of an experiment and are not useful here. given some $\alpha\in\mathbb{R}$ are of the type
$f_{\alpha}\left(\lambda\right)=\left|\lambda\right|^{\alpha}.$ (4.48)
Going forward, we will assume $\alpha\neq 0$ and continuity of these $f$-functions, which makes the $p$-functions continuous as well:
$p\left(\lambda\mathbf{a}\right)=\left|\lambda\right|^{\alpha}p\left(\mathbf{a}\right).$ (4.49)
Functions of this type are called positively homogeneous of degree $\alpha$.
We find for the edge cases
$p\left(\lambda\mathbf{1}\right)=0$ $\quad\longrightarrow\quad\lambda=0,$ (4.50)
$p\left(\lambda\mathbf{1}\right)=1$ $\quad\longrightarrow\quad\lambda=1.$ (4.51)
4.8 Scalar field
Real numbers $\mathbb{R}$ are an algebraic field. To prove consistency with the scalar representation of symmetric paths, we now trace all field axioms for $\left\langle\mathbb{R},\oplus,\odot\right\rangle$ from the model where possible and otherwise show that they are not in conflict:
  • Coarsening of two symmetric paths always results in another symmetric path, and the result is, therefore, again a scalar. This also applies to chains of symmetric paths. The operation $\oplus$ is closed.
  • Additive associativity and commutativity of $\oplus$ are satisfied from general coarsening, and additive inverses always exist.
  • Impossible paths have amplitude $\mathbf{0}$, which is the additive identity, $\mathbf{a}\oplus\mathbf{0}=\mathbf{0}\oplus\mathbf{a}=\mathbf{a}$.
  • Chains of symmetric paths are also represented as scalars, and the operation $\odot$ is closed.
  • Multiplicative associativity is satisfied from associativity of chaining $\odot$.
  • Multiplicative commutativity is required by postulate from real-number representation. It constrains the algebra but is not in conflict with the model.
  • Paths that are trivial in their nonredundant form are also trivially symmetric. Chaining with a trivial path is redundant and corresponds to applying a scalar multiplicative unit, $1\mathbf{a}\equiv\mathbf{a}$.
  • Multiplicative inverses of $\mathbf{a}$ exist for $\mathbf{a}\neq\mathbf{0}$ as long as $p\left(\mathbf{a}\right)\neq 0$. Because $p\left(\lambda\mathbf{1}\right)=0$ can only be true when $\lambda=0$, inverses exist for all nonzero scalars.
  • Distributivity of multiplication over addition corresponds to distributivity of chaining over coarsening in general.
With this, amplitudes representing chains of palindromic sequences satisfy all field axioms when identifying field multiplication with chaining and field addition with coarsening. We point out that most of the field axioms trace from model properties of symmetric paths, except for commutativity of multiplication, which is postulated.
Choosing the real numbers rules out other possible realizations of an algebraic field. For example, requiring scalars to be ordered and one-dimensional rules out the two-dimensional complex numbers $\mathbb{C}$ and specific field extensions (e.g. by supplying the rationals $\mathbb{Q}$ with a second dimension to basis $\sqrt{2}$). In one dimension, fixing the reals excludes fields over a set with smaller cardinality (e.g., the rationals $\mathbb{Q}$), fields with characteristic not 0 (e.g., finite fields, or the so-called “nimbers” which are infinite), and fields over a set that is congruent to the reals but with a different valuation (such as the $p$-adic numbers). These interesting opportunities are advertised for exploration in follow-on work.
4.9 Vector space
We now recognize the amplitude space $\left\langle\mathbb{A},\oplus,\odot\right\rangle$ to be a vector space. In the most general case, a vector space $V$ over a field $K$ is a set with two operations, vector addition $\oplus:V\times V\rightarrow V$ and scalar multiplication with a vector $\otimes:K\times V\rightarrow V$, such that for all vectors $\mathbf{a},\mathbf{b},\mathbf{c}\in V$ and scalars $\lambda,\mu\in K$, there is
  1. 1.
    $\left(\mathbf{a}\oplus\mathbf{b}\right)\oplus\mathbf{c}=\mathbf{a}\oplus\left(\mathbf{b}\oplus\mathbf{c}\right)$,
  2. 2.
    $\mathbf{a}\oplus\mathbf{b}=\mathbf{b}\oplus\mathbf{a}$,
  3. 3.
    there exists an element $\mathbf{0}\in V$ such that $\mathbf{0}\oplus\mathbf{a}=\mathbf{a}\oplus\mathbf{0}=\mathbf{a}$,
  4. 4.
    there exists an element $\left(-\mathbf{a}\right)\in V$ such that $\mathbf{a}\oplus\left(-\mathbf{a}\right)=\mathbf{0}$,
  5. 5.
    $\lambda\otimes\left(\mu\otimes\mathbf{a}\right)=\left(\lambda\mu\right)\otimes\mathbf{a}$,
  6. 6.
    $1\otimes\mathbf{a}=\mathbf{a}$,
  7. 7.
    $\lambda\otimes\left(\mathbf{a}\oplus\mathbf{b}\right)=\lambda\otimes\mathbf{a}\oplus\lambda\otimes\mathbf{b}$,
  8. 8.
    $\left(\lambda+\mu\right)\otimes\mathbf{a}=\lambda\otimes\mathbf{a}\oplus\mu\otimes\mathbf{a}$.
From these general axioms, we identify elements $\mathbf{a},\mathbf{b},\mathbf{c}\in V$ with amplitudes $\mathbf{a},\mathbf{b},\mathbf{c}\in\mathbb{A}$, the field $K$ with the reals $\mathbb{R}$, and scalars $\lambda,\mu\in K$ with amplitudes $\lambda\mathbf{1},\mu\mathbf{1}\in\mathbb{A}$. Axioms “1.” and “2.” then follow directly from coarsening. The additive identity $\mathbf{0}\in\mathbb{A}$ represents impossible paths (“3.”), and amplitude addition is invertible (“4.”). Scalars $\lambda\equiv\lambda\mathbf{1}$ mix with general amplitudes exactly as in “5.” through “8.” when identifying scalar multiplication with amplitude multiplication, $\otimes\equiv\odot$.
There are known theorems for faithful representation of a vector space. A linearly independent set of vectors that spans the entire vector space is called a basis of the vector space. A vector space generally has many bases, all of which have the same cardinality. This is the dimension of the vector space.
After fixing such a basis $\mathbf{e}_{V}$ of an $n$-dimensional vector space $V$ to the $n$-tuple $\mathbf{e}_{V}:=\left(\mathbf{e}_{0},\ldots,\mathbf{e}_{n-1}\right)$, $\mathbf{e}_{r}\in V$ for $r=0,\ldots,n-1$, every vector $\mathbf{a}\in V$ can be written uniquely as a linear combination of these basis elements with scalars $a_{r}\in K$:
$\mathbf{a}=:\sum_{r=0}^{n-1}a_{r}\mathbf{e}_{r}.$ (4.52)
The scalars $a_{r}$ are called coordinates (or coefficients) of the vector $\mathbf{a}$, which can then be represented as tuple to the corresponding basis elements in $\mathbf{e}_{V}$:
$\mathbf{a}=\left(a_{0},\ldots,a_{n-1}\right).$ (4.53)
4.10 Algebra over a field
The configuration space of amplitudes is a real vector space that is further equipped with a bilinear product, i.e., a multiplication $\odot$ that is both homogeneous
$\left(\lambda\mathbf{a}\right)\odot\left(\mu\mathbf{b}\right)=\lambda\mu\left(\mathbf{a}\odot\mathbf{b}\right),$ (4.54)
and distributes over addition $\oplus$,
$\mathbf{a}\odot\left(\mathbf{b}\oplus\mathbf{c}\right)$ $=\mathbf{a}\odot\mathbf{b}\oplus\mathbf{a}\odot\mathbf{c},$ (4.55)
$\left(\mathbf{b}\oplus\mathbf{c}\right)\odot\mathbf{a}$ $=\mathbf{b}\odot\mathbf{a}\oplus\mathbf{c}\odot\mathbf{a}.$ (4.56)
This characterizes the amplitude algebra $\mathbb{A}$ as a conventional algebra over a field, here over the reals $\mathbb{R}$.
Given a basis $\mathbf{e}_{V}:=\left(\mathbf{e}_{0},\ldots,\mathbf{e}_{n-1}\right)$ of an $n$-dimensional vector space $V$, structure constants $f_{ijk}\in K$ with $i,j,k=0,\ldots,n-1$ allow to explicitly represent the product of $\mathbb{A}$, by specifying the product of every pair of basis vectors $\mathbf{e}_{i},\mathbf{e}_{j}$ as a linear combination:
$\mathbf{e}_{i}\odot\mathbf{e}_{j}:=\bigoplus_{k=0}^{n-1}f_{ijk}\mathbf{e}_{k}.$ (4.57)
Because every vector $\mathbf{a}\in V$ can be written as a linear combination of the $\mathbf{e}_{i}$, this product extends to all vectors in the vector space.
Every product from an algebra over a field can be uniquely represented by structure constants (after fixing a basis). Conversely, any product represented by structure constants is from an algebra over a field.
4.11 $p$-functions as signomials
The positively homogeneous $p$-functions of degree $\alpha$ over vectors $\mathbf{a}=\left(a_{0},\ldots,a_{n-1}\right)$ are now expressed as generalized polynomials over the vector coefficients. Referred to as a signomials, these are sums of arbitrarily many (generalized) monomials, indexed $s=1,2,3,\ldots$ and weighted by some $c_{s}\in\mathbb{R}$. These monomials are products made from the coefficients of $\mathbf{a}$, forced nonnegative, and taken to some real powers $\beta\left(s,r\right)\in\mathbb{R}_{\geq 0}$ that add up to $\alpha$:
$p\left(\mathbf{a}\right):=\sum_{\forall s}c_{s}\left(\prod_{r=0}^{n-1}\left|a_{r}\right|^{\beta\left(s,r\right)}\right),\textrm{ with }\sum_{r=0}^{n-1}\beta\left(s,r\right)=\alpha\textrm{ for all }s$ (4.58)
(using $0^{0}:=1$ by convention). Since the exponents in all monomials add up to $\alpha$, this can be termed a homogeneous signomial [^12] For contemporary use of term, and signomials in general, see e.g. [DM2022]. of degree $\alpha$.
The reader can readily verify that (4.58) solves the homogeneity equation for the $p$-functions [^13] The authors were not able to find completeness or uniqueness theorems for signomials towards solving the homogeneity equation in the current literature, or likewise, for solutions of the multiplicative Cauchy equation (Cauchy’s “power equation”) $p\left(\mathbf{x}\odot\mathbf{y}\right)=p\left(\mathbf{x}\right)p\left(\mathbf{y}\right)$ for general vector products $\mathbf{x}\odot\mathbf{y}$. Works like [AD1989] give solutions for matrix products in the most general case. It may be interesting to ask whether, for example, the expression $\left(a_{0}^{2}+a_{1}^{2}\right)^{\pi}$ for $a_{0},a_{1}\in\mathbb{R}$ can be written as an infinite sum of products of the $a_{0},a_{1}$ taken to exponents that add up to $2\pi$, i.e., $\left(a_{0}^{2}+a_{1}^{2}\right)^{\pi}=\left|a_{0}\right|^{2\pi}+\left|a_{1}\right|^{2\pi}+\ldots+c_{s}\left|a_{0}\right|^{\beta\left(s,0\right)}\left|a_{1}\right|^{\beta\left(s,1\right)}+\ldots$ with $\beta\left(s,0\right)+\beta\left(s,1\right)=2\pi$ for all $s$. Since signomials of the form (4.58) will prove to be sufficient to specify the $p$-function in this paper, we leave these follow-on questions for the interested reader., $p\left(\lambda\mathbf{a}\right)=\left|\lambda\right|^{\alpha}p\left(\mathbf{a}\right)$.
5 Born rule
Connecting to canonical physics, we now show that $p$-function is quadratic in amplitudes, and interpret it as a generalized Born rule from Quantum Mechanics. Using well-known axioms from mathematics, we constrain allowable amplitude algebras to the small, distinct set of associative composition algebras over the reals. This includes the complex numbers, concluding our argument towards the suitability of the model and methodology in this paper.
5.1 Quadratic form $Q\left(\mathbf{a}\right)$
The product $\mathbf{a}\odot\overline{\mathbf{a}}$ is self-adjoint for any $\mathbf{a}\in\mathbb{A}$, and therefore, always corresponds to a scalar, $\mathbf{a}\odot\overline{\mathbf{a}}=\lambda\mathbf{1}$, $\lambda\in\mathbb{R}$, per equation (4.43). We now define a map $Q:\mathbb{A}\rightarrow\mathbb{R}$ as:
$Q\left(\mathbf{a}\right)\mathbf{1}$ $:=\mathbf{a}\odot\overline{\mathbf{a}}=\lambda\mathbf{1}.$ (5.1)
Given $\mu\in\mathbb{R}$, there is:
$Q\left(\mu\mathbf{a}\right)\mathbf{1}=\left(\mu\mathbf{1}\odot\mathbf{a}\right)\odot\overline{\left(\mu\mathbf{1}\odot\mathbf{a}\right)}=\mu^{2}Q\left(\mathbf{a}\right)\mathbf{1}.$
Because $Q$ is positively homogeneous of degree $2$, it can be expressed as a homogeneous signomial over the coefficients of $\mathbf{a}$ similar to the $p$-functions. Furthermore, since it is a product from an algebra over a field, expressible through summation of structure constants, it is, in fact, a homogeneous polynomial of degree $2$. This makes $Q$ a quadratic form.
From associativity of amplitudes and commutativity of scalars follows
$Q\left(\mathbf{a}\right)\mathbf{1}\odot\mathbf{a}$ $=\left(\mathbf{a}\odot\overline{\mathbf{a}}\right)\odot\mathbf{a}=\mathbf{a}\odot\left(\overline{\mathbf{a}}\odot\mathbf{a}\right)=\mathbf{a}\odot Q\left(\overline{\mathbf{a}}\right)\mathbf{1}=Q\left(\overline{\mathbf{a}}\right)\mathbf{1}\odot\mathbf{a},$ (5.2)
$Q\left(\mathbf{a}\right)$ $=Q\left(\overline{\mathbf{a}}\right).$ (5.3)
Also, $Q$ is multiplicative since for any $\mathbf{b}\in\mathbb{A}$ with $\mathbf{b}\odot\overline{\mathbf{b}}=\mu\mathbf{1}$, $\mu\in\mathbb{R}$ there is
$Q\left(\mathbf{a}\odot\mathbf{b}\right)\mathbf{1}$ $=\mathbf{a}\odot\mathbf{b}\odot\overline{\mathbf{a}\odot\mathbf{b}}=\mathbf{a}\odot\mathbf{b}\odot\overline{\mathbf{b}}\odot\overline{\mathbf{a}}$
$=\mathbf{a}\odot\left(\mathbf{b}\odot\overline{\mathbf{b}}\right)\odot\overline{\mathbf{a}}=\mu\mathbf{a}\odot\overline{\mathbf{a}}=\lambda\mu\mathbf{1}$ (5.4)
$=Q\left(\mathbf{a}\right)Q\left(\mathbf{b}\right)\mathbf{1}.$
5.2 Quadratic probabilities
To show the behavior of the $p$-functions under scaling, we first look at a symmetric experiment $^{\circledcirc}\negthinspace\mathcal{M}=\left[\mathbf{N},\mathbf{M},\mathbf{N}\right]$, where measurement $\mathbf{N}$ has at least one atomic detector $N$, and $\mathbf{M}$ has $s$ atomic detectors, $\mathbf{M}=\left\{M^{1},M^{2},\ldots M^{s}\right\}=\left\{\left\{m^{1}\right\},\left\{m^{2}\right\},\ldots\left\{m^{s}\right\}\right\}$. Fixing $N$ as source and target, the possible symmetric paths $^{\circledcirc}\negthinspace A^{r}\in\mathcal{S}_{{}^{\circledcirc}\negthinspace\mathcal{M}}$ with $r=1\ldots s$ are:
$^{\circledcirc}\negthinspace A^{r}:=\left[N,M^{r},N\right].$ (5.5)
When coarse-graining over all paths, $^{\cup}A:=\bigvee_{r=1}^{s}\,{}^{\circledcirc}\negthinspace A^{r}=\left[N,\,^{\cup}M,N\right]$, the measurement
$^{\cup}\mathbf{M}=\left\{{}^{\cup}M\right\}=\left\{\bigcup_{r=1}^{s}M^{r}\right\}=\left\{\left\{m^{1},m^{2},\ldots m^{s}\right\}\right\}$ (5.6)
is certain and may be removed from $^{\cup}A$ without changing its probability. The remaining trivial path $\left[N,N\right]$ has unit amplitude, such that the fully coarse-grained path must have unit probability,
$P\left({}^{\cup}A\right)=P\left(\bigvee_{r=1}^{s}\,{}^{\circledcirc}\negthinspace A^{r}\right)=1.$ (5.7)
All paths $^{\circledcirc}\negthinspace A^{r}$ are symmetric and therefore represented by scalars, making $\phi\left({}^{\cup}A\right)=\bigoplus_{r=1}^{s}\phi\left(\,{}^{\circledcirc}\negthinspace A^{r}\right)=\lambda\mathbf{1}$ a scalar as well. $P\left({}^{\cup}A\right)=p\left(\phi\left({}^{\cup}A\right)\right)=p\left(\lambda\mathbf{1}\right)=1$ infers $\lambda=1$ per (4.51), and fully coarsening results in the unit amplitude:
$\phi\left({}^{\cup}A\right)=\mathbf{1}.$ (5.8)
Paths $^{\circledcirc}\negthinspace A^{r}$ can be factorized into pairs of adjoint paths because all detectors $M^{r}:=\left\{m^{r}\right\}$ are atomic:
$^{\circledcirc}\negthinspace A^{r}$ $\sim\left[N,M^{r}\right]\cdot\left[M^{r},N\right].$ (5.9)
With amplitudes $\mathbf{a}^{r}:=\phi\left(\left[N,M^{r}\right]\right)$ we have therefore:
$\phi\left(A^{r}\right)$ $=\phi\left(\left[N,M^{r}\right]\right)\odot\phi\left(\left[M^{r},N\right]\right)=\mathbf{a}^{r}\odot\overline{\mathbf{a}^{r}},$ (5.10)
$\phi\left({}^{\cup}A\right)$ $=\bigoplus_{r=1}^{s}\left(\mathbf{a}^{r}\odot\overline{\mathbf{a}^{r}}\right)=\mathbf{1}.$ (5.11)
Using the quadratic form $Q$, this becomes:
$\bigoplus_{r=1}^{s}\left(\mathbf{a}^{r}\odot\overline{\mathbf{a}^{r}}\right)$ $=\bigoplus_{r=1}^{s}Q\left(\mathbf{a}^{r}\right)\mathbf{1}=\mathbf{1},$ (5.12)
$\sum_{r=1}^{s}Q\left(\mathbf{a}^{r}\right)$ $=1.$ (5.13)
Conversely, the probability of independently observing any of the paths $\left[N,M^{r}\right]$ must also be $1$ per section 3.9,
$\sum_{r=1}^{s}P\left(\left[N,M^{r}\right]\right)=\sum_{r=1}^{s}p\left(\mathbf{a}^{r}\right)=1.$ (5.14)
The sum of quadratic forms as in (5.13) is again a quadratic form, and the space of signomials of the same degree is closed under summation likewise (5.14). For both these sums to always add up to $1$ over the same coefficients from all $\mathbf{a}_{r}$, the exponential behavior of all summands from the argument components must be identical. This identifies $\alpha=2$ and the $p$-functions with $Q$,
$p$ $\equiv Q,$ (5.15)
$p\left(\nu\mathbf{1}\right)$ $=\nu^{2},$ (5.16)
$p\left(\nu\mathbf{a}\right)$ $=\nu^{2}p\left(\mathbf{a}\right).$ (5.17)
Because the $p$-functions model probability, when traced from valid path operations in the model, we have effectively shown that probabilities are quadratic in transition amplitudes between measurements. This can be interpreted as a generalized Born rule, in the sense of Feynman’s rules as in [GKS2010].
5.3 $p$-functions are a nondegenerate quadratic form
To fully utilize our configuration space of the amplitude algebra, we want the $p\left(\mathbf{a}\right)$ to be nondegenerate: They must require $n$ independent parameters in the $n$-dimensional vector space of $\mathbb{A}$ and cannot be reduced to a function over a lower-dimensional subspace of $\mathbb{A}$.
This can be formulated in a coordinate-independent way by using the associated bilinear form $B:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$,
$B\left(\mathbf{a},\mathbf{b}\right)\mathbf{1}:=\left(p\left(\mathbf{a}\oplus\mathbf{b}\right)-p\left(\mathbf{a}\right)-p\left(\mathbf{b}\right)\right)\mathbf{1}=\mathbf{a}\odot\overline{\mathbf{b}}\oplus\mathbf{b}\odot\overline{\mathbf{a}}.$ (5.18)
Degeneracy of $B$ and, by extension, $p$, then corresponds to existence of a nonzero $\mathbf{b}_{\bot}\in\mathbb{A}$ for which $B\left(\mathbf{a},\mathbf{b}_{\bot}\right)=0$ for all $\mathbf{a}$. Such a nondegenerate $B$ allows to conclude for any given $\mathbf{c}\in\mathbb{A}$
$B\left(\mathbf{a},\mathbf{c}\right)=0\textrm{ for all }\mathbf{a}\in\mathbb{A}\rightarrow\mathbf{c}=\mathbf{0}.$ (5.19)
To show equivalence of such nondegeneracy with our requirement for $n$ linear independent arguments modeling $p$ in an $n$-dimensional vector space, we show that a degenerate $p$ can always be modeled with fewer than $n$ independent parameters; and conversely that a nondegenerate $p$ requires $n$ parameters.
First, we assume that $p$ would be degenerate, and write $\mathbf{b}_{\bot}\in\mathbb{A}$ for a nonzero amplitude that has $B\left(\mathbf{a},\mathbf{b}_{\bot}\right)=0$ for all $\mathbf{a}$. In that case equation (5.18) becomes $p\left(\mathbf{a}\oplus\mathbf{b}_{\bot}\right)=p\left(\mathbf{a}\right)+p\left(\mathbf{b}_{\bot}\right)$ and $p$ is an additive map for every $\mathbf{a}$. This is trivially satisfied if $p$ is constant 0. For nonconstant $p$, Cauchy’s functional equation would require the continuous $p$ to be linear in $\mathbb{R}$. Since we already established $p$ to be quadratic in general, this can only be the case for general $\mathbf{a}$ if $p\left(\mathbf{b}_{\bot}\right)=0$. This makes $p\left(\mathbf{a}\oplus\mu\mathbf{b}_{\bot}\right)=p\left(\mathbf{a}\right)+\mu^{2}p\left(\mathbf{b}_{\bot}\right)=p\left(\mathbf{a}\right)$ for any $\mu\in\mathbb{R}$, and $p$ cannot depend on the one-dimensional space spanned by $\mu\mathbf{b}_{\bot}$. That is, degenerate $p$ can always be reduced to a function that excludes the parameters where $p\left(\mu\mathbf{b}_{\bot}\right)=0$ and, therefore, fully understood using fewer than $n$ independent parameters.
Conversely, now assuming $p$ would be nondegenerate but depending on fewer than $n$ independent coefficients of its argument. It would mean that a coordinate basis exists in which the 0-coefficient of any vector does not contribute to the value of $p$. In such a basis, choose a vector $\mathbf{b}^{\prime}:=\left(b_{0}^{\prime},0,\ldots,0\right)$ that differs from the zero-vector $\left(0,\ldots,0\right)$ only by $b_{0}^{\prime}\neq 0$. Since $p\left(\mathbf{b}^{\prime}\right)$ does not depend on the 0-component of its argument, there is always $p\left(\mathbf{b}^{\prime}\right)=p\left(\mathbf{0}\right)=0$, and likewise $p\left(\mathbf{a}\oplus\mathbf{b}^{\prime}\right)=p\left(\mathbf{a}\right)$. With this, $B\left(\mathbf{a},\mathbf{b}^{\prime}\right)=p\left(\mathbf{a}\oplus\mathbf{b}^{\prime}\right)-p\left(\mathbf{a}\right)-p\left(\mathbf{b}^{\prime}\right)=0$ for every $\mathbf{a}$, which contradicts the requirement for $p$ being nondegenerate.
5.4 Composition algebras
Per [Jacobson1958], a unital algebra $\mathbb{A}$ over the reals [^14] The treatment in [Jacobson1958] is more general in that it allows not just the real number field but any field $K$ of characteristic not 2. We will not go into details here since we have already chosen $K=\mathbb{R}$ per postulate. is a composition algebra if it admits a nondegenerate quadratic form $Q:\mathbb{A}\rightarrow\mathbb{R}$ with product composition. This means that there exists a multiplicative unit $\mathbf{1}\in\mathbb{A}$, and for every $\mathbf{a},\mathbf{b}\in\mathbb{A}$, $\lambda\in\mathbb{R}$ there is:
$Q\left(\mathbf{a}\odot\mathbf{b}\right)$ $=Q\left(\mathbf{a}\right)Q\left(\mathbf{b}\right),$ (5.20)
$Q\left(\lambda\mathbf{a}\right)$ $=\lambda^{2}Q\left(\mathbf{a}\right),$ (5.21)
$B\left(\mathbf{a},\mathbf{b}\right)$ $:=Q\left(\mathbf{a}+\mathbf{b}\right)-Q\left(\mathbf{a}\right)-Q\left(\mathbf{b}\right),$ (5.22)
$B\left(\mathbf{a},\mathbf{c}\right)=0\textrm{ for all }\mathbf{a}\in\mathbb{A}$ $\rightarrow\mathbf{c}=\mathbf{0}.$ (5.23)
We have traced all these axioms from the model, concluding that any allowable amplitude algebra must be a composition algebra.
Only a few distinct algebras over the real numbers admit these properties. Next to the field algebra $\mathbb{R}$ itself are the complex $\mathbb{C}$, quaternion $\mathbb{H}$, and octonion $\mathbb{O}$ algebras over vector spaces $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, and $\mathbb{R}^{8}$, respectively, with their split forms $\mathbb{C}^{\prime}$, $\mathbb{H}^{\prime}$, and $\mathbb{O}^{\prime}$. Because we require multiplication to be associative, the nonassociative octonions $\mathbb{O}$ and split-octonions $\mathbb{O}^{\prime}$ are ruled out here [^15] For composition algebras in general, and these possibilities over the reals in particular, see, for example [PM:CompAlg, WP:CompAlg]..
Our $p$-function on amplitudes is the well-known quadratic form $Q$ on these algebras, which may either be positive-definite (for $\mathbb{R}$, $\mathbb{C}$, $\mathbb{H}$) or have split-signature (for $\mathbb{C}^{\prime}$, $\mathbb{H}^{\prime}$). For $\mathbb{R}$, $\mathbb{C}$ and $\mathbb{H}$ it is the Euclidean norm of $\mathbf{a}$ in one, two, and four dimensions, respectively:
$p_{\mathbb{R}}\left(\mathbf{a}\right)=p_{\mathbb{R}}\left(\left(a_{0}\right)\right)$ $=a_{0}^{2},$ (5.24)
$p_{\mathbb{C}}\left(\mathbf{a}\right)=p_{\mathbb{C}}\left(\left(a_{0},a_{1}\right)\right)$ $=a_{0}^{2}+a_{1}^{2},$ (5.25)
$p_{\mathbb{H}}\left(\mathbf{a}\right)=p_{\mathbb{H}}\left(\left(a_{0},a_{1},a_{2},a_{3}\right)\right)$ $=a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}.$ (5.26)
Conversely, for the split-algebras $\mathbb{C}^{\prime}$ and $\mathbb{H}^{\prime}$ it is an isotropic quadratic form that is hyperbolic [^16] Given a suitable labeling of basis elements. We will not further go into the details of split-algebras in this paper, but point to popular introductions, e.g., as in [WP:SplitQuaternion].:
$p_{\mathbb{C}^{\prime}}\left(\mathbf{a}\right)=p_{\mathbb{C}^{\prime}}\left(\left(a_{0},a_{1}\right)\right)$ $=a_{0}^{2}-a_{1}^{2},$ (5.27)
$p_{\mathbb{H}^{\prime}}\left(\mathbf{a}\right)=p_{\mathbb{H}^{\prime}}\left(\left(a_{0},a_{1},a_{2},a_{3}\right)\right)$ $=\left(a_{0}^{2}+a_{1}^{2}\right)-\left(a_{2}^{2}+a_{3}^{2}\right).$ (5.28)
We point out that the dimensionality of allowable amplitude algebras, $\dim\mathbb{A}\in\left\{1,2,4\right\}$, has not been prescribed or postulated by us anywhere but follows solely from the nature of composition algebras.
Expressions in conventional Quantum Mechanics typically are over complex numbers and often use matrix formulations. While quaternions or split-algebras may not appear familiar at first sight, they are effectively already part of the standard toolset of a physicist: The familiar Pauli matrices are typically represented as $2\times 2$ complex matrices and can be used to construct the imaginary parts of a quaternion basis. The algebra of $2\times 2$ real matrices is, in fact, a split-quaternion algebra[WP:SplitQuaternion].
With this, we have demonstrated the utility of complex numbers in our model and related composition algebras already used in canonical physics today. This supports our claim that the model can be a building block towards a faithful, complete reconstruction of the laws of Quantum Mechanics.
5.5 Alternative axioms towards composition algebras
Next to the axioms used above, Nathan Jacobson gave two more alternative sets of axioms, each leading to the same composition algebras over the reals. In order to strengthen our argument for internal consistency of our axioms and choices, we now discuss these as well [^17] While the power of axiomatic, deductive reasoning could be paraphrased as “proven once, valid everywhere,” choice representations may carry doubt of their universality. This could be put pointedly maybe as “chosen once, to be proven everywhere.” In practice, the situation is not nearly as dire, as choices at some point become axioms by their very nature. Nevertheless, until a complete axiomatic description of Quantum Mechanics is posed, it is prudent to check for internal consistency between our current choices and the more foundational axioms..
Also in [Jacobson1958], a unital alternative algebra $\mathbb{A}$ over $\mathbb{R}$ is a composition algebra if it admits an involution $\mathbf{a}\rightarrow\overline{\mathbf{a}}$, such that both product and sum of any $\mathbf{a}\in\mathbb{A}$ with its involution $\overline{\mathbf{a}}$ is a real form. This means that there exists a multiplicative unit $\mathbf{1}\in\mathbb{A}$, and for every $\mathbf{a},\mathbf{b}\in\mathbb{A}$, there is:
$\mathbf{a}\odot\left(\mathbf{a}\odot\mathbf{b}\right)$ $=\left(\mathbf{a}\odot\mathbf{a}\right)\odot\mathbf{b},$ (5.29)
$\mathbf{b}\odot\left(\mathbf{a}\odot\mathbf{a}\right)$ $=\left(\mathbf{b}\odot\mathbf{a}\right)\odot\mathbf{a},$ (5.30)
$\overline{\overline{\mathbf{a}}}$ $=\mathbf{a},$ (5.31)
$\mathbf{a}\odot\overline{\mathbf{a}}$ $=Q\left(\mathbf{a}\right)\mathbf{1}\textrm{ with }Q:\mathbb{A}\rightarrow\mathbb{R},$ (5.32)
$\mathbf{a}\oplus\overline{\mathbf{a}}$ $=T\left(\mathbf{a}\right)\mathbf{1}\textrm{ with }T:\mathbb{A}\rightarrow\mathbb{R}.$ (5.33)
We readily recognize all axioms but the last one from our work here. Interestingly, these axioms do not mention the product composition property of the $Q$ or existence of a quadratic form. Instead, properties of the involution are required, which in turn are absent from the earlier axiom set.
Algebraically, we can immediately verify that the $T$-function (the “trace”) is indeed real from (5.18):
$B\left(\mathbf{a},\mathbf{1}\right)\mathbf{1}=\mathbf{a}\oplus\overline{\mathbf{a}}.$ (5.34)
This is, therefore, consistent with our model. However, this amplitude expression can never represent an allowable path operation,
$\phi\left(A\overset{??}{\vee}\overline{A}\right)=\mathbf{a}\oplus\overline{\mathbf{a}}.$ (5.35)
Looking at $A\overset{??}{\vee}\overline{A}$ by itself, $A$ would have to differ from $\overline{A}$ by exactly one measurement result to permit coarsening, which is impossible. Conversely, in an expression
$A\overset{??}{\vee}\left(\overline{A}\vee B\right)$ (5.36)
with some path $B$, the cardinality of one of the measurement results in $A$ would have to be larger than the corresponding one in $\overline{A}$, which is a contradiction. The trace function is an example of an algebraic property of the representation space that is formally always true but can never be traced from the model; and is, therefore, not in conflict with the model, either.
We can however trace a third set of axioms towards composition algebras, given in [Jacobson1981]. Only looking at the real algebra case, an alternative algebra $\mathbb{A}$ over $\mathbb{R}$ is a composition algebra if it is free from absolute zero divisors and admits an involution $\mathbf{a}\rightarrow\overline{\mathbf{a}}$ that is an anti-isomorphism, where the product of any $\mathbf{a}\in\mathbb{A}$ with its involution $\overline{\mathbf{a}}$ is a real form. This means that for every $\mathbf{a},\mathbf{b},\mathbf{c}\in\mathbb{A}$ there is:
$\mathbf{a}\odot\left(\mathbf{a}\odot\mathbf{b}\right)$ $=\left(\mathbf{a}\odot\mathbf{a}\right)\odot\mathbf{b},$ (5.37)
$\mathbf{b}\odot\left(\mathbf{a}\odot\mathbf{a}\right)$ $=\left(\mathbf{b}\odot\mathbf{a}\right)\odot\mathbf{a},$ (5.38)
$\overline{\overline{\mathbf{a}}}$ $=\mathbf{a},$ (5.39)
$\mathbf{a}\odot\overline{\mathbf{a}}$ $=Q\left(\mathbf{a}\right)\mathbf{1}\textrm{ with }Q:\mathbb{A}\rightarrow\mathbb{R},$ (5.40)
$\overline{\mathbf{a}\odot\mathbf{b}}$ $=\overline{\mathbf{b}}\odot\overline{\mathbf{a}},$ (5.41)
$\left(\mathbf{c}\odot\mathbf{a}\right)\odot\mathbf{c}=\mathbf{0}\textrm{ with }\mathbf{a}\neq\mathbf{0}$ $\rightarrow\mathbf{c}=\mathbf{0}.$ (5.42)
Again, all but the last of the axioms are familiar and were previously shown to trace from the model. As for the last item, for $\left(\mathbf{c}\odot\mathbf{a}\right)$ to result in a zero amplitude for any nonzero $\mathbf{a}$, the $\mathbf{c}$ must represent an impossible path and be $\mathbf{0}$, making the entire expression $\left(\mathbf{c}\odot\mathbf{a}\right)\odot\mathbf{c}=\mathbf{0}$. Where $\left(\mathbf{c}\odot\mathbf{a}\right)$ would be nonzero, conversely, chaining another $\odot\mathbf{c}$ could make that expression $\mathbf{0}$ only if $\mathbf{c}=\mathbf{0}$, which is a contradiction. Amplitude algebra is, therefore, free from absolute zero divisors, and all axioms are traced from our model.
5.6 A note on split-algebras
Split-complex numbers $\mathbb{C}^{\prime}$ and split-quaternions $\mathbb{H}^{\prime}$ contain subspaces where the $p$-function is 0 for some nonzero amplitudes. In $\mathbb{C}^{\prime}$ with $p_{\mathbb{C}^{\prime}}\left(\mathbf{a}\right)=a_{0}^{2}-a_{1}^{2}$, these are the nonzero diagonals in the plane, $a_{0}=\pm a_{1}$. In $\mathbb{H}^{\prime}$ with $p_{\mathbb{H}^{\prime}}\left(\mathbf{a}\right)=\left(a_{0}^{2}+a_{1}^{2}\right)-\left(a_{2}^{2}+a_{3}^{2}\right)$ these are the hyperplanes with $a_{0}^{2}+a_{1}^{2}=a_{2}^{2}+a_{3}^{2}$. As section 4.4 points out, such algebraic subspaces must never represent allowable paths in the model algebra. Furthermore, $\mathbb{C}^{\prime}$-spaces $\left|a_{0}\right|<\left|a_{1}\right|$ and $\mathbb{H}^{\prime}$-spaces $a_{0}^{2}+a_{1}^{2}<a_{2}^{2}+a_{3}^{2}$ have a negative $p$-function, which can never occur in $\mathbb{C}$ or $\mathbb{H}$.
While this may appear undesirable at first, the situation is essentially no different, e.g., in the complex numbers $\mathbb{C}$ where numbers outside the unit circle have a $p$-function value greater than $1$, which would be nonsensical when interpreted as probability.
To be consistent with our model, we do have to satisfy one requirement: As long as the $p$-function represents paths with nonzero probability, multiplication must be invertible (per section 4.6.2). This is indeed the case: Given any $\mathbf{a}\in\mathbb{A}\in\left\{\mathbb{R},\mathbb{C},\mathbb{C}^{\prime},\mathbb{H},\mathbb{H}^{\prime}\right\}$, multiplicative inverses $\mathbf{a}^{-1}\in\mathbb{A}$ are well-defined for nonzero $p\left(\mathbf{a}\right)$:
$\mathbf{a}^{-1}=\frac{\overline{\mathbf{a}}}{p_{\mathbb{A}}\left(\mathbf{a}\right)}\quad\left(p_{\mathbb{A}}\left(\mathbf{a}\right)\neq 0\right).$ (5.43)
This is, therefore, consistent with the required existence of multiplicative inverses for amplitudes representing allowable paths.
Compared to the split-algebras with their hyperbolic $p$-function, the circular geometry in $\mathbb{C}$ and $\mathbb{H}$ may appear simpler or more symmetric when separating formally defined abstract amplitudes from those with $p\in\left[0,1\right]$ that may represent concrete paths in the model. However, this does not exclude split-algebras from utility towards reconstructing Quantum Mechanics: The experiment ultimately verifies nature’s geometry, regardless of how simple or symmetric we may find it.
6 Summary and outlook
6.1 What has and has not been accomplished
In this paper, we explored a model that was first advertised in [GKS2010] as a building block in the Quantum Reconstruction Program, which included reconstructing core algebraic properties brought initially forward by Richard P. Feynman. Addressing open questions posed in [Koepl2023gks], we confirmed that it is possible to represent probabilities from the model by amplitudes on complex numbers. Going beyond the original work, we did so coordinate-independently without requiring a two-dimensional configuration space a priori. Furthermore, we traced algebraic unit elements, scalar multiplication, and most field axioms for scalars from the model. Demonstrating that scalars form an algebraic field, including invertibility of addition and multiplication, allowed us to conclude that the configuration space of amplitudes is indeed a vector space [^18] We believe this to be a significant innovation, also with thanks to John A. Shuster for pointing out a counterexample early on: If one were to define multiplication $\odot$ and addition $\oplus$ between a pair of two-dimensional vectors $\mathbf{a}$ and $\mathbf{b}$ in terms of familiar complex number operations, $\mathbf{a}\odot\mathbf{b}:=\exp\left(\ln\mathbf{a}\ln\mathbf{b}\right)$, $\mathbf{a}\oplus\mathbf{b}:=\mathbf{a}\mathbf{b}$, then such a system would satisfy many axioms of a conventional algebra over a field, including distributivity of multiplication over addition; however, scalar multiplication would not be homogeneous since for example $\left(\lambda\mathbf{a}\right)\odot\mathbf{b}=\exp\left(\ln\left(\lambda\mathbf{a}\right)\ln\mathbf{b}\right)=\exp\left(\ln\lambda\ln\mathbf{b}\right)\left(\mathbf{a}\odot\mathbf{b}\right)\neq\lambda\left(\mathbf{a}\odot\mathbf{b}\right)$. It would be consistent with axioms stated in the original work but not support the conclusion. In general, algebraic constructs that result from dropping homogeneity of vector multiplication, $\left(\lambda\mathbf{a}\right)\odot\left(\mu\mathbf{b}\right)=\lambda\mu\left(\mathbf{a}\odot\mathbf{b}\right)$, while maintaining distributivity of multiplication over addition, have been investigated under the “commutative hyperoperation” label and proposed as early as [Bennett1915].. We showed that probabilities are quadratic in amplitudes without needing to know the dimensionality of the vector space. All along, we followed a program that aimed at providing clarity of thought at the interface of mathematics and physics: We separated axioms from choices in an attempt to differentiate what properties of the model derive from foundational assumptions, in contrast to properties that trace to choices made by an outside experimenter in order to provide answers for specific questions.
Curiously, the purely mathematical nature of composition algebras over the reals prescribes the dimensionality of allowable amplitude algebras. We found no grounds for excluding the two-dimensional split-complex numbers from permissible amplitude algebras, as well as the four-dimensional quaternions and split-quaternions. Quaternions do not occur in the original work due to the fixed requirement of a two-dimensional configuration space. To rule out the split-algebras in two and four dimensions, there would have to be some triangle inequality constraint for $\sqrt{p}$ (e.g. [Baez2002]), possibly from convexity axioms on probabilities; or alternatively, required existence of a nondegenerate trilinear form in the amplitude algebra (triality; also [Baez2002]). We did not find a corresponding structure in the model from which such could be traced.
While we weakened or eliminated several postulates from the original work under investigation, the following algebraic properties still have to be required ad-hoc, i.e., without tracing to the model or the question for probabilities: commutativity of scalar multiplication and continuity of the $p$-function. These seemingly small postulates have far-reaching consequences: Allowing scalars to be noncommutative would generalize the space of amplitudes from being a vector space to merely a module. It would require clarification on how to generalize the quadratic form $Q$ and its behavior under involution of the argument (5.3), and even more so, it would entirely leave the scope of Jacobson’s axioms classifying composition algebras. Conversely, without continuity of the $p$-function, its exponential behavior (4.49) and signomial representation of the $p$-functions (4.58) could not be deduced. Not knowing whether $p\left(\lambda\mathbf{1}\right)=0$ infers $\lambda=0$ (4.50) would leave a gap in satisfying the field axiom for multiplicative inverses for scalars. Not knowing whether $p\left(\lambda\mathbf{1}\right)=1$ infers $\lambda=1$ (4.51) would prevent the conclusion that fully coarsening scalars must result in the unit amplitude (5.8). All of these are required to show that the $p$-functions are indeed a quadratic form.
As a final mathematical call-out, we motivated real-number representation of the scalar field from a simplicity argument in section 4.8 to keep the mathematical space of amplitudes as similar as possible to the space of probabilities they ought to explain. It is nevertheless a relatively strong choice. As hinted at with examples in that section, weakening this choice could take the work in different directions.
From the physics point of view, we hope that our improved clarity in treating this elementary building block of quantum reality provides insight into existing follow-on work that considers composite systems [Goyal2014] and distinguishability [Goyal2015] and allows the model to be incorporated into new or existing work that reconstructs spacetime, the light cone, and the dynamical interaction of quantum systems [^19] Prominent graph-based approaches that address these questions include Causal Set theory, which offers powerful insights into proposed structures of spacetime, and the intriguing spin networks of Loop Quantum Gravity. For recent review articles, see e.g. [Rovelli2012, Surya2019, AB2021, Surya2025]..
6.2 Histories matter for amplitudes?
When we asked for probabilities of measurement outcomes in section (3), we required that any experiment be independent of the system’s history prior to the first measurement. This concept of closure was part of our question; hence, we required it from the representation from which we expect our answer. In future work, we could subtly modify our question, to still require that histories of a system do not matter for probabilities of a subsequent measurement; however, there may exist further degrees of freedom in the amplitude algebra that do not contribute to future probabilities. These degrees of freedom may depend on a system’s history and, with that, no longer be subject to closure as defined in this paper. In a way, this weaker form of closure would resemble an anthropocentric stance, where properties of nature that are conceptually out of reach for a human observer cannot be prepared for in an experiment, are driven by the system’s past, and consequently drive behavior of that system that likewise remains out of human reach.
Regardless of whether the current paper’s concept of (strong) closure or the proposed weak closure would be assumed when formulating questions to the system, the resulting answers for probabilities would remain the same. That does not mean, however, that postulating weak closure would therefore be inconsequential. To sketch a window of opportunity for this ansatz, its outcome could be that the degrees of freedom inaccessible to humans contribute to measurement outcomes in a way that would make certain aspects of the outcome follow seemingly random patterns. Those patterns could be irreconcilable with classical probability yet follow well-defined probability distributions. Nature, by supposition, would answer a question for probabilities by fitting its inaccessible inner workings into the template provided by that very question. This sketch traces observed properties from canonical Quantum Mechanics, of course.
Under weak closure, it would not be allowable anymore to take an amplitude of an existing path and prepend a new measurement to it, as this could amount to altering the past of the existing path. Because histories now matter, we would only be allowed to model further experiments by appending measurements to existing ones. In formulas, starting with two measurement sequences $\mathcal{M}^{\mathrm{A}}$ and $\mathcal{M}^{\mathrm{B}}$ that are chained, $\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}$, we would only be allowed to chain a third sequence $\mathcal{M}^{\mathrm{C}}$ to the side of the expression that represents the future (say, to the right) because chaining it to the side that represents the past (the left) would correspond to altering history under weak closure. Expressions like $\left(\mathcal{M}^{\mathrm{A}}\cdot\mathcal{M}^{\mathrm{B}}\right)\cdot\mathcal{M}^{\mathrm{C}}$ would remain allowable, but expressions from the mere regrouping of the order of brackets, $\mathcal{M}^{\mathrm{A}}\overset{??}{\cdot}\left(\mathcal{M}^{\mathrm{B}}\cdot\mathcal{M}^{\mathrm{C}}\right)$, could not be traced from the model anymore.
Algebraically, this permits algebras, where multiplication is not necessarily associative. The reader readily recognizes the nonassociative octonions $\mathbb{O}$ and split-octonions $\mathbb{O}^{\prime}$ to become permissible [Okubo1995, Baez2002, ConwaySmithD2003] since associativity of multiplication was the only property that ruled out these real composition algebras in this paper. Just as for the complexes and quaternions, their $p$-function models probability, which does remain invariant when modifying product association. This, therefore, still satisfies the weak closure requirement for probabilities.
While algebraic nonassociativity may feel inconvenient at first, it is essentially no different than noncommutativity in the quaternions $\mathbb{H}$, which likewise does not affect the $p$-function of the product [^20] Or noncommutativity in matrix multiplication, which does not affect the determinant.. Even more so, building an algebra from one-sided multiplication only and letting products act on a separate element $\psi$ of that algebra multiplicatively forms an associative algebra by definition: It simply forces all brackets a certain way. Using such one-sided multiplication algebras built from octonions with real coefficients, $\mathbb{O}$, with complex coefficients $\mathbb{C}\otimes\mathbb{O}$, quaternionic coefficients $\mathbb{H}\otimes\mathbb{O}$, and iteratively further “tensoring” these to form algebras like $\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ and larger, all form Clifford algebras; and that separate element $\psi$ these act on multiplicatively hence is identified as a spinor [^21] By definition, a spinor can be understood as an element on which a Clifford algebra acts.. This construction has been explored in physics for use in modeling symmetries in the Standard Model of particles since the late 20th century [Dixon1994, Dixon2014]. Significant recent progress towards a deeper, more complete understanding of the various algebraic aspects, possible realizations, degrees of freedom, consequences, and constraints [Furey2016, FuH2022-OneGen, FuH2022-SymmBr, Furey2023rmap1] led to active research in the field. Curiously, these endeavors use the same algebras that would become permissible in our model here when allowing weak closure for answering an experimenter’s question.
Prescribing the order of multiplication also means that all algebraic properties must now only be satisfied in one-sided multiplication, namely, for the side from which it is allowed to multiply a priori. Sometimes called “(left-/right-)multiplication algebras,”[Schafer1966] there are initial discoveries of normed multiplication algebras that permit a linear, homogeneous multiplication from one side but are nondistributive and nonhomogeneous when multiplying from the other [SmithJDH1995, ConwaySmithD2003, SmithW2004]. A complete mathematical classification is subject to ongoing investigation.
In nonassociative multiplication algebras, there is no concept of inserting unit amplitudes $\mathbf{1}$ through chaining at different places in a measurement sequence. Such an operation would not be allowed, as you may multiply only from one side. With that, the existence of a multiplicative unit amplitude $\mathbf{1}$ in the algebra becomes optional, permitting so-called nonunital algebras. This is distinct from the existence of a unit scalar, $1$, which can always be modeled into any sequence through redundancy from duplicated measurements. In our work, scalars are separate from amplitudes by definition, which is consistent with the new freedom. Nonunital algebras over the reals include the eight-dimensional Okubo algebra [Okubo1995, SmithVojtech2022]. In the reals, normed nonunital algebras with product composition still permit triality [Elduque1996, Elduque2000, Elduque2022, KO2015, KO2016]. This key property of algebras that would become permissible in our model, namely, of generally not associative or unital multiplication algebras with triality [^22] Classification of these algebras is an open question in mathematics as well., is likewise explored in contemporary algebraic physics research [Hughes2016, Koepl2023aut, MCZ2023, FuH2024].
6.3 Do all histories matter?
Assuming that all histories matter when calculating amplitudes, an earlier graph-based model was captured in [Furey2016] (section 2) that postulates for nodes from all elements in the lower set of a partially ordered set (from the “poset ideal”) to contribute to an algebraic expression. When asking for an observation, that expression will be evaluated; otherwise, when left untouched by an observer, the expression will keep accumulating nodes indefinitely as the system evolves. This extreme concept of “all histories matter” supports an intuition towards amplitude representation in the form of algebraic ideals, i.e., algebraic subspaces that absorb general amplitudes into their respective subspace under multiplication yet are closed under amplitude addition. Under some conservation assumptions, suitable algebras in which these ideals possess symmetries that resemble the ones observed in the Standard Model of particles in physics are found in tensored division algebras over the reals, $\left(\mathbb{R}\otimes\right)\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ [Furey2010] – yet again.
The occurrence of related algebraic structures in different contexts, now even tracing from different model assumptions, leads one to wonder about a common provenance. How could our current “histories don’t matter” postulate be reconcilable with an “all histories matter” axiom? Closing on an inspirational note, when weakening closure from “histories don’t matter” to include “… for probabilities only,” to become the anthropocentric consequence of an observer approaching nature with the very question for probabilities, it is conceivable that our model here describes the behavior of quantum nature under observation [^23] Quoting [Goyal2012]: Quantum Mechanics may be “[not] a description of reality in itself [but] a description of reality as experienced by an agent”.; whereas “all histories matter” approaches like superdeterminism[AHHP2024] describe the inner workings of nature when left unobserved [^24] For a sketch of an agency-free model of Quantum Field Theory, on algebras permissible when extending our current work, see e.g. [Furey2023osmu]..
Acknowledgments
We would like to extend our sincere thanks to Nichol Furey, Bernd Henschenmacher, John Huerta, and John A. Shuster for their helpful discussion, pointers, corrections, and mentoring, as well as to the organizers and attendees of the “Mile High Conferences on Nonassociative Mathematics” (2009, 2013, and 2017) at the University of Denver.